### v1

# Med-R1: Reinforcement Learning for Generalizable Medical Reasoning in Vision-Language Models

Abstract
Vision-language models (VLMs) have made significant progress in reasoning within natural scenes, yet their potential in medical imaging remains largely underexplored. Medical reasoning tasks, which require robust image analysis and the generation of well-justified answers, present unique challenges due to the inherent complexity of medical images. Transparency and trustworthiness are critical not only for gaining clinicians’ confidence but also for meeting stringent regulatory requirements. To address these challenges, we propose Med-R1, a novel framework that investigates whether reinforcement learning (RL) can enhance the generalizability and trustworthiness of VLMs in medical reasoning. Building on the recently introduced DeepSeek strategy, we adopt Group Relative Policy Optimization (GRPO) for RL, which encourages models to explore reasoning paths guided by reward signals. Unlike supervised fine-tuning (SFT), which often overfits to training data and struggles with generalization, RL enables models to develop more robust and diverse reasoning capabilities. We comprehensively evaluate Med-R1 across eight distinct medical imaging modalities: Computed Tomography (CT), Magnetic Resonance Imaging (MRI), Ultrasound, Dermoscopy, Fundus Photography, Optical Coherence Tomography (OCT), Microscopy Images, and X-ray Imaging. Compared to the base model, Qwen2-VL-2B, Med-R1 achieves a 29.94% improvement in average accuracy across these modalities and even outperforms Qwen2-VL-72B—a model with 36 times more parameters. To assess model’s generalization abilities, we further test on five different question types: modality recognition, anatomy identification, disease diagnosis, lesion grading, and biological attribute analysis. Med-R1 also demonstrates superior cross-task generalization, outperforming Qwen2-VL-2B by 32.06% and Qwen2-VL-72B in question-type generalization accuracy. These results highlight that RL not only enhances medical reasoning capabilities but also enables parameter-efficient models to exceed the performance of significantly larger counterparts. Combined with its interpretable reasoning outputs, Med-R1 establishes a new potential direction for developing generalizable, trustworthy, and clinically deployable medical VLMs.
1 Introduction
Vision-language models (VLMs) have achieved remarkable progress in reasoning over natural images (e.g., GPT-4o [16], Gemini-1.5 [31], and Qwen-VL [2]), demonstrating impressive capabilities in tasks such as visual question answering (VQA) and multimodal dialogue. These advancements have been driven by large-scale vision-language pretraining and supervised fine-tuning (SFT), which enable models to associate visual content with textual information. However, applying VLMs to medical imaging reasoning remains a significant challenge. Unlike natural images, medical data demand precise interpretation, requiring VLMs to generate not only final conclusions but also step-by-step reasoning processes that align with clinical logic. For instance, diagnosing a lung nodule in a CT scan necessitates localizing lesions, analyzing morphological features, and integrating patient history—a reasoning chain that must be both accurate and interpretable to gain clinicians’ trust. Furthermore, the diversity of medical imaging modalities (e.g., CT, MRI, microscopy) and task types (e.g., diagnosis, lesion grading) poses stringent demands on model generalizability. These challenges raise a critical question: How can we equip VLMs with robust reasoning abilities that generalize across medical domains while maintaining clinical trustworthiness?
In this paper, we identify that the limitations of current medical VLMs primarily stem from the inherent drawbacks of Supervised Fine-Tuning (SFT) [6, 1]. While SFT has been widely adopted to adapt foundation models to medical imaging [23, 34, 18, 5, 21], it suffers from two fundamental issues that hinder clinical applicability. First, SFT inherently biases models toward memorizing task-specific shortcuts rather than learning generalizable reasoning. By directly aligning model outputs with final answers (e.g., diagnostic labels), SFT encourages overfitting to superficial patterns in training data. Second, the scarcity of high-quality Chain-of-Thought (CoT) annotations severely limits the effectiveness of SFT in medical reasoning. Unlike general-domain tasks, where large-scale CoT datasets can be crowdsourced, medical reasoning requires domain-specific logical structuring (e.g., systematically ruling out differential diagnoses before confirming malignancy). However, curating such CoT datasets is prohibitively expensive, as it demands meticulous annotation by experienced medical professionals to ensure diagnostic validity and clinical coherence. As a result, existing SFT-based medical VLMs [23, 34] lack access to high-quality CoT data, leading to shallow and non-clinically grounded reasoning. These models frequently produce "black-box" predictions, failing to generate interpretable rationales or adapt to out-of-domain tasks (e.g., from radiology to pathology). This lack of transparency and robustness poses a significant challenge to clinical adoption, where explainability and reliability are paramount.
To address these challenges, we propose Med-R1, a framework leveraging Reinforcement Learning (RL) [27] to enhance both the generalizability and trustworthiness of medical VLMs. Unlike SFT, RL guides models to explore diverse reasoning pathways through reward signals, encouraging them to develop intrinsic logical steps rather than relying on annotated CoT rationales. This approach mitigates overfitting and shortcut learning, as demonstrated by Chu et al. [9], who showed that RL-trained vision-language models significantly outperform SFT in cross-domain generalization. Building on recent advancements in RL optimization, we adopt Group Relative Policy Optimization (GRPO) [28], a novel strategy that eliminates the need for complex value models required by traditional methods like Proximal Policy Optimization (PPO) [27]. GRPO incorporates rule-based rewards derived from medical guidelines (e.g., diagnostic criteria for tumors) and group-relative comparisons to stabilize policy updates. For example, rewards can be designed to prioritize reasoning steps that align with radiological decision trees, ensuring clinically plausible rationales. This dual mechanism not only reduces computational overhead but also enforces adherence to domain-specific constraints, making GRPO uniquely suited for medical reasoning tasks where transparency and scalability are paramount.
Figure 1: Overview of the Evaluation Framework: Eight Medical Imaging Modalities and Five Medical Vision Question Answering Tasks. We evaluate Med-R1 across eight distinct medical imaging modalities—Computed Tomography (CT), Magnetic Resonance Imaging (MRI), Ultrasound, Dermoscopy, Fundus Photography, Optical Coherence Tomography (OCT), Microscopy Images, and X-ray Imaging—as well as five medical vision question answering tasks: anatomy identification, disease diagnosis, lesion grading, modality recognition, and biological attribute analysis. Example images and corresponding clinical questions illustrate the diversity of medical data and the challenges in developing generalizable vision-language models for automated medical reasoning.
As shown in Figure 1, to rigorously validate the effectiveness of Med-R1, we conduct a comprehensive evaluation spanning eight clinically critical imaging modalities: Computed Tomography (CT), Magnetic Resonance Imaging (MRI), Ultrasound, Dermoscopy, Fundus Photography, Optical Coherence Tomography (OCT), Microscopy Images, and X-ray Imaging. These modalities were selected to represent the spectrum of medical imaging challenges, from macroscopic anatomical visualization (CT/MRI) to cellular-level analysis (Microscopy) and functional assessments (OCT). Across these domains, Med-R1 (2B parameters) achieves an average accuracy of 69.91%, representing a 29.94% absolute improvement over its foundation model Qwen2-VL-2B. Notably (Table 2), Med-R1 even surpasses Qwen2-VL-72B—a 72B-parameter model with 36× larger scale, demonstrating that parameter efficiency need not compromise performance when leveraging RL-driven reasoning. Beyond evaluating performance across imaging modalities, we also examine Med-R1’s generalizability across different medical question types. We assess the model’s capability in five key question categories: modality recognition, anatomy identification, disease diagnosis, lesion grading, and biological attribute analysis. Med-R1 exhibits superior cross-task generalization, outperforming Qwen2-VL-2B by 32.06% in question-type generalization accuracy, and even surpassing Qwen2-VL-72B in this aspect (Table 4). These findings emphasize the potential of RL-based training in improving both reasoning depth and task adaptability in medical vision-language models (VLMs).
Med-R1 is a generalizable medical vision-language model (VLM) that generates clinically interpretable reasoning explanations for medical visual question answering (VQA) across eight imaging modalities (Figure 2). We summarized our contribution as follows:
  1. 1.
    Generalizable Medical VLM with Multi-Modal Explicit Reasoning: We propose Med-R1, the first vision-language model supporting eight medical imaging modalities (CT, MRI, Ultrasound, etc.), capable of generating step-by-step reasoning explanations without task-specific retraining. Unlike prior SFT-based models limited to single modalities and black-box answers, Med-R1 achieves cross-domain generalization through RL-driven exploration.
  2. 2.
    Emergent Reasoning via Clinically Constrained RL: We adopt Group Relative Policy Optimization (GRPO), a RL strategy that eliminates the need for explicit reasoning supervision. GRPO integrates rule-based rewards (e.g., enforcing adherence to diagnostic guidelines) and group-relative comparisons, enabling models to learn clinically plausible reasoning paths from final-answer labels alone.
  3. 3.
    Robust Generalization with Efficiency: Med-R1 outperforms foundation model by 29.94% and SFT baselines by 15.84% in cross-modality accuracy. We also enhance 32.06% on foundation modal and 11.25% than SFT baselines in cross-task generalization. Moreover, Med-R1 surpasses other larger models in the general domain or medical domain including Qwen2-VL-72B, and MedVInT-7B, demonstrating scalability and trustworthiness for real-world deployment.
2 Related Works
General VLMs and Medical VLMs. Vision-language models (VLMs) have significantly advanced multimodal understanding by aligning visual and textual representations through large-scale pretraining. General-purpose VLMs, such as CLIP [26] and BLIP-2 [19], excel in natural image-text tasks but encounter challenges when applied to domain-specific medical reasoning. Recent efforts have sought to adapt these models to medical imaging through supervised fine-tuning (SFT) on domain-specific datasets. For instance, LLaVA [22] introduces instruction tuning for general-purpose vision-language tasks, achieving strong performance on natural image benchmarks. Building on this, LLaVA-Med [18] fine-tunes LLaVA using radiology reports, demonstrating improved performance on medical visual question answering (VQA) tasks. Similarly, Med-Flamingo [23] extends Flamingo by incorporating domain-specific pretraining for medical VQA. However, these SFT-based approaches often suffer from overfitting to narrow medical corpora, limiting their generalizability across diverse imaging modalities and clinical tasks. Our work addresses this gap by introducing reinforcement learning (RL) for generalization, enabling robust performance without modality-specific retraining.
Reinforcement Learning for Post-Training. Reinforcement learning [30, 8] has emerged as a powerful tool for aligning large language models (LLMs) with human preferences through reward signals. Approaches like RLHF [24] and RLAIF [4] leverage human or AI-generated feedback to refine model behavior. In the vision-language domain, RL has been applied to improve VQA accuracy [9] and reduce hallucination [28]. However, existing methods rely on complex reward models or extensive human annotation, limiting scalability in data-scarce domains like medicine. Our Group Relative Policy Optimization (GRPO) addresses these limitations by incorporating rule-based rewards derived from clinical guidelines and group-relative comparisons, enabling parameter-efficient adaptation to diverse medical tasks.
Medical Reasoning and Interpretability Recent work emphasizes the importance of interpretable reasoning in medical AI, with approaches like Chain-of-Thought (CoT) prompting [33] and program-guided reasoning [6]. However, acquiring such annotations in the medical domain is particularly challenging due to the need for expert involvement [20], high annotation costs, and the complexity of medical reasoning. The RL-based approach enables emergent reasoning without explicit supervision, addressing this challenge.
Figure 2: Medical VQA examples of Med-R1 The left panel demonstrates a modality recognition task, where the model correctly identifies the presence of breast tissue in a mammogram (X-ray). The right panel illustrates a disease diagnosis task, where Med-R1 accurately detects urolithiasis in an abdominal CT scan. In both cases, the model provides interpretable reasoning (“Think”), explaining its decision-making process before selecting the final answer.
3 Method
We adopt RL to encourage multimodal reasoning and generalization in medical domains. Recent studies have demonstrated that RL can incentivize emergent logical reasoning and generalization in multimodal tasks such as mathematical reasoning [15] and visual navigation [9]. Building on these insights, we extend its application to the medical domain and systematically assess its effectiveness in this context. Specifically, we leverage the popular RL-based post-training method GRPO [29] to train a large base MLLM [3] across 8 different modalities for medical reasoning and compare it with zero-shot and SFT performance of popular existing VLMs. We introduce the details of the algorithm, our reward design, and data structure below.
3.1 Group Relative Policy Optimization (GRPO)
Overview.
RL-based algorithms such as PPO [27] and GRPO [28] belong to a family of fine-tuning and alignment strategies explicitly designed to enhance models’ reasoning capacities. Unlike supervised fine-tuning, which directly optimizes maximum likelihood, these RL-based methods instead optimize the policy gradient using reward signals, encouraging reasoning by exploring a much larger solution space. GRPO is closely related to PPO but differs in two key aspects: first, GRPO estimates the advantage using group-based estimation rather than a value function; second, it employs a set of fixed rules as the reward signal instead of a learned reward model. These optimizations make GRPO 50% more resource- and computation-efficient than PPO [28].
Definition.
Formally, let $P(Q)$ denote the question set used for training, where $q$ is a sampled question in the current iteration. Let $\pi_{\theta_{\text{old}}}$ and $\pi_{\theta_{\text{new}}}$ denote the old policy and current (new) policy, respectively, where $o$ is a complete response sampled from a policy. Let $\pi_{\theta_{\text{ref}}}$ denote the reference policy, which in practice is the frozen base MLLM. Let $G$ be the number of responses sampled per question in each iteration. The GRPO objective is given by:
$\mathcal{J}_{\text{GRPO}}(\theta)$ $=\mathbb{E}_{q\sim P(Q),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O|q)}$ (1)
$\frac{1}{G}\sum_{i=1}^{G}\Bigg{(}\min\Bigg{(}\frac{\pi_{\theta_{\text{new}}}(o_{i}|q)}{\pi_{\theta_{\text{old}}}(o_{i}|q)}A_{i},\text{clip}\Bigg{(}\frac{\pi_{\theta_{\text{new}}}(o_{i}|q)}{\pi_{\theta_{\text{old}}}(o_{i}|q)},1-\epsilon,1+\epsilon\Bigg{)}A_{i}\Bigg{)}-\beta\mathbb{D}_{KL}\Big{(}\pi_{\theta_{\text{new}}}\Big{\|}\pi_{\text{ref}}\Big{)}\Bigg{)}$
where $\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\text{old}}}(o_{i}|q)}$ is the policy ratio, and $A_{i}$ is the estimated advantage. The KL divergence term [17] regularizes the policy update, ensuring that $\pi_{\theta}$ does not deviate excessively from the reference model $\pi_{\theta_{\text{ref}}}$. Unlike PPO, which uses a critic model to estimate the advantage $A_{i}$ for a single response $o$, GRPO estimates the relative advantage by sampling a group of responses $\{o_{i}\}_{i=1}^{G}$ and normalizing their rewards within the group to compute a relative advantage [28, 13]. Each reward is calculated based on rules without using a regard model. We detail the reward design below.
Reward design.
We follow [28] and use two types of reward: format and accuracy. To begin with, we prompt the model to explicitly output its thinking process in the “<think>…</think>" tag and the final answer in the “<answer>…</answer>" tag. The format reward is designed to check if the aforementioned tags are present in the final response. A reward score of 1 will be given if they exist and are correct. This helps the model to organize its thoughts and answer in a structured format for the ease of reading. The accuracy reward is a rule-based reward that checks if the actually answer matches with the ground truth. Similarly, a reward score of 1 is given when the results match. In practice, the ground truths are letter options “A, B, C, D" for multiple choice questions, and we treat all responses with correct letter options as the leading word as correct (“A…").
Table 1: Cross-Modality Generalization Performance of Med-R1 with RL Post-Training. Accuracy (%) across eight medical imaging modalities, where rows indicate training modalities and columns test modalities. Darker cell shades indicate higher accuracy for corresponding training-test pairs. The "Overall" column/row reports average performance per training and test domain.
RL fine-tuned VLM
Train Test CT MRI X-Ray US Der FP OCT Micro Overall
CT 98.3 74.43 80.61 51.93 64.17 65.39 70.52 66.13 71.44
MRI 62.33 99.29 73.75 53.18 75.04 64.57 75.59 66.31 71.26
X-Ray 79.67 68.29 95.85 49.32 70.75 76.87 70.28 67.75 72.35
US 55.88 63.96 70.53 99.86 64.78 63.02 63.92 71.53 69.19
Der 53.81 67.36 73.31 49.80 95.10 70.22 64.03 67.93 67.71
FP 54.98 65.95 75.79 49.08 71.36 92.44 65.33 66.40 67.67
OCT 64.48 70.80 77.09 52.65 72.89 73.41 99.17 66.85 72.17
Micro 60.97 63.25 74.92 52.56 64.55 63.66 66.86 93.51 67.54
Overall 66.30 71.67 77.73 57.31 72.33 71.20 71.96 70.80 69.91
  • CT - Computed Tomography; MRI - Magnetic Resonance Imaging; US - Ultrasound.
  • Der - Dermoscopy; FP - Fundus Photography; OCT - Optical Coherence Tomography
  • Micro - Microscopy Images; X-Ray - X-Ray Imaging
4 Experiment & Results
In this section, we introduce our experimental setup, implementation details, and results. We present a preliminary analysis over the results and discuss more in LABEL:sec:_discussion.
4.1 Setup
Datasets. We adopt the VQA data from the open-access part of the OmniMedVQA benchmark [14], which consists of a total of $82,059$ images and $88,996$ vision question answering (VQA) pairs. OmniMedVQA includes VQA pairs from eight imaging modalities: CT ($15,808$), MRI ($31,877$), X-Ray ($7,916$), Ultrasound ($10,991$), Dermoscopy ($6,679$), Fundus ($5,398$), OCT ($,64$ 646), and Microscopy ($5,680$). It is also categorized into five VQA question types, including Anatomy Identification ($16,448$), Disease Diagnosis ($55,387$), Lesion Grading ($2,098$), Modality Recognition ($11,565$), and Other Biological Attributes ($3,498$). We split the dataset into training and test sets following an 80-20 ratio for each setting.
Implementation Details Training is conducted on HGX H100 [7] server with 2×H100 GPUs (80GB VRAM) using PyTorch [25] and FlashAttention-2 [11] for optimized efficiency. We initialize from Qwen2-VL-2B-Instruct [2] with full parameter tuning, employing per-GPU batch size 1 (effective batch size 4 via 2-step gradient accumulation) and bfloat16 mixed precision. Input sequences combine visual embeddings from 328×328 resolution images (max 401k pixels) with textual prompts truncated to 1,024 tokens. The GRPO policy generates four candidate rationales per sample, with a sampling temperature of $\tau=0.7$. Each training configuration is run for one epoch.
Task setting. We evaluate our approach in two distinct generalization settings using the OmniMedVQA dataset [14]: cross-modality generalization and cross-task generalization.
  • Cross-modality generalization: We train our model on a single modality at a time (out of the 8 available modalities) and evaluate its performance on the remaining 7 unseen modalities.
  • Cross-task generalization: We identify 5 distinct task types within the dataset and adopt the same train-test partitioning strategy as in the cross-modality setting, training on one task and evaluating on the remaining four.
Baseline Methods & Evaluation Metric. Our evaluation framework integrates three classes of baselines: general VLMs including BLIP-2 [19], InstructBLIP [10], and the Qwen2-VL series (2B/7B/72B) [2, 32]; medical VLMs pretrained on domain-specific corpora (Med-Flamingo [23], MedVInT [34], LLaVA-Med [18]); and supervised fine-tuned (SFT) variants of Qwen2-VL-2B.
We evaluate model performance using VQA choice accuracy, the standard metric for medical VQA, where models select answers from clinically validated options. Given a medical image $I$, a question $Q$, and $K$ candidate answers $\{A_{k}\}_{k=1}^{K}$ (one correct and $K-1$ distractors), accuracy is computed as:
$\text{Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left(\hat{y}_{i}=y_{i}\right)$ (2)
where $N$ is the total number of test cases, $\hat{y}_{i}$ is the model’s predicted answer index, $y_{i}$ is the ground truth index, and $\mathbb{I}$ is the indicator function, which returns 1 if the predicted answer matches the ground truth and 0 otherwise.
Table 2: Performance Comparison of VLMs on Eight Modality: Performance comparison across eight medical imaging modalities demonstrates that our GRPO-finetuned model surpasses zero-shot general-purpose VLMs, domain-specific medical VLMs, and conventional supervised fine-tuning approaches, achieving superior accuracy while maintaining scalability for clinical deployment. Meanwhile, in each column, the best and second-best performances are marked in red and blue. Naming details are as follows: CT - Computed Tomography; MRI - Magnetic Resonance Imaging; US - Ultrasound; Der - Dermoscopy; FP - Fundus Photography; OCT - Optical Coherence Tomography; Micro - Microscopy Images; X-Ray - X-Ray Imaging.
Methods Modality CT MRI X-Ray US Der FP OCT Micro Overall
Zero-Shot VLM
[…]
$\text{InstructBLIP}^{\dagger}$ [10] 28.72 33.15 61.04 41.25 62.22 50.31 42.59 46.29 45.70
[…]
Qwen2-VL-2B [2] 45.10 38.57 39.32 30.86 35.83 43.17 35.14 36.85 38.11
Qwen2-VL-7B [2] 61.46 45.77 64.27 36.01 49.08 59.84 59.32 61.08 54.60
Qwen2-VL-72B [32] 67.97 69.39 77.21 51.39 65.31 72.58 72.76 67.83 68.05
Zero-Shot Medical VLM
[…]
$\text{MedVInT}^{\dagger}$ [34] 40.74 43.10 55.10 41.26 29.11 31.84 23.26 32.00 37.05
[…]
fine-tuned VLM
Qwen2-VL-2B (SFT) 51.74 52.83 65.57 47.65 51.91 52.26 53.99 56.58 54.07
Ours-2B (GRPO) 66.30 71.67 77.73 57.31 72.33 71.20 71.96 70.80 69.91
4.2 Cross-Modality Generalization
We comprehensively evaluate Med-R1’s adaptability across 8 medical imaging modalities, including Computed Tomography (CT), Magnetic Resonance Imaging (MRI), Ultrasound, Dermoscopy, Fundus Photography, Optical Coherence Tomography (OCT), Microscopy Images, and X-ray Imaging. Our experiments focus on two key aspects: (1) cross-modal generalization, where the model is trained on one modality and tested on another, and (2) comparative performance against other popular vision-language models (VLMs) and medical-specific VLMs evaluated using zero-shot and SFT.
Results on generalization. To evaluate Med-R1’s cross-modality generalization, we measure its accuracy across eight distinct medical imaging modalities (Table 1). The Overall row and column summarize the model’s average performance across training and test domains, providing insights into its generalization ability. Med-R1 achieves a strong overall accuracy of 69.91%, demonstrating its ability to generalize across diverse medical imaging modalities. Notably, models trained on CT, MRI, and X-Ray exhibit the highest generalization capability, with overall scores of 71.44%, 71.26%, and 72.35%, respectively. In contrast, models trained on Fundus Photography (FP) and Microscopy (Micro) images show lower generalization, with 67.67% and 67.54% overall accuracy, indicating that certain modality-specific features (e.g., texture-based imaging in US and Micro) may not transfer as effectively to other domains. Importantly, the overall test accuracy of 69.91% highlights Med-R1’s ability to perform well across unseen imaging modalities, despite being trained on a single domain at a time. This result underscores the effectiveness of reinforcement learning (RL) in enhancing cross-modality transfer, allowing the model to maintain robust performance without requiring extensive retraining for each medical imaging modality.
Comparison to zero-shot and SFT evaluations with other VLMs. As demonstrated in Table 2, Med-R1 demonstrates its superiority across all eight medical imaging modalities while maintaining exceptional parameter efficiency. For zero-shot results, each cell denotes the zero-shot evaluation accuracy of the model on the particular modality. For the fine-tuned VLM results (last two rows), each cell represents the overall accuracy, reflecting the average generalization performance of the given modality when evaluated using models that were separately fine-tuned on each of the eight training modalities. Against general-purpose VLMs, our 2B-parameter model achieves 69.91% overall accuracy, surpassing the 72B-parameter Qwen2-VL by 1.86%—a notable result given the 36×36$\times$36 × parameter disparity. This advantage amplifies in critical diagnostic tasks: Med-R1 attains 71.67% accuracy in MRI compared to Qwen2-VL-72B’s 69.39%, and achieves 72.33% versus 65.31% in dermoscopy, challenging the prevailing scale-equals-performance paradigm. The limitations of specialized medical VLMs become evident through Med-Flamingo’s 30.38% average accuracy, which Med-R1 outperforms by 39.53%. This stark contrast underscores the ineffectiveness of narrow medical pretraining compared to our RL-driven adaptation strategy. When compared to supervised fine-tuning approaches, GRPO delivers 15.84% accuracy gains over SFT-tuned Qwen2-VL-2B (69.91% vs. 54.07%), with particularly significant improvements in CT interpretation (66.30% vs. 51.74%) and OCT analysis (71.96% vs. 53.99%).
Table 3: Cross-Task Generalization of Med-R1: Performance is evaluated across five clinical reasoning task types (rows: training tasks, columns: test tasks), with darker cell shading indicating stronger generalization. Results demonstrate that domain-specific training (e.g., Disease Diagnosis) preserves in-task expertise while maintaining adaptability to unseen tasks, particularly for modality-agnostic skills like Modality Recognition.
Train Test Anatomy Disease Lesion Modality Other Overall
Identification Diagnosis Grading Recognition Attributes
Anatomy Identification 96.06 59.07 55.51 98.62 74.86 76.83
Disease Diagnosis 54.16 98.25 73.85 97.87 84.09 81.64
Lesion Grading 52.30 57.72 86.24 97.32 74.29 73.57
Modality Recognition 55.30 55.43 56.88 99.46 70.88 67.59
Other Attributes 56.6 59.93 56.88 97.91 96.59 73.58
Overall 62.88 66.08 65.87 98.24 80.14 74.64
4.3 Cross-Task Generalization
We also evaluate Med-R1’s generalization across 5 important clinical tasks [14]: Anatomy Identification, Disease Diagnosis, Lesion Grading, Modality Recognition, and Other Attributes. Similar to subsection 4.2, we focus our evaluation on two aspects: cross-modality generalization and comparison against SFT and zero-shot with other VLMs.
Results on generalization. As shown in Table 3, models trained on “disease diagnosis" data achieve the best generalization, with 81.64% overall accuracy. This suggests that disease diagnosis encompasses diverse feature representations that transfer well across tasks, likely due to its reliance on both anatomical and pathological cues. In contrast, models trained on “modality recognition" exhibit strong generalization in task-agnostic settings (98.24% in the “modality recognition" column), indicating that learning modality distinctions aids in extracting transferable image features. However, training on “lesion grading" leads to high in-task performance (86.24%) but relatively lower transferability, implying that this task captures more specialized features that do not generalize as effectively. These results highlight the trade-off between specialization and adaptability, emphasizing the importance of task selection when designing models for broad medical applications.
Comparison to zero-shot and SFT evaluations with other VLMs. Table 4 shows the comparison results with other popular VLMs evaluated with zero-shot and SFT. For zero-shot results, each cell denotes the zero-shot evaluation accuracy of the model on the particular task. For the fine-tuned VLM results (last two rows), each cell represents the overall accuracy, reflecting the average generalization performance of the given task when evaluated using models that were separately fine-tuned on each of the five training tasks. First of all, the results clearly show that Med-R1 outperforms all other popular VLMs’ zero-shot generalization. Remarkably, Med-R1 even outperforms Qwen2-VL-72B (74.64% vs. 72.58%), a model with 70 billion more parameters. More importantly, this suggests that RL can effectively elevate small models with moderate capacity, opening doors for many real-world applications where resource is a constraint. In contrast, the average generalization with the identical base model trained with SFT is merely 63.39%, 11.25% below Med-R1, further demonstrating the strong generalization capability of Med-R1.
Table 4: Performance Comparison of VLMs on Five Medical VQA Tasks: GRPO Fine-Tuning Outperforms Zero-Shot and SFT Baselines Across Diverse Reasoning Tasks. Performance is evaluated on five clinical reasoning types (columns) across three model categories: general-purpose VLMs (zero-shot), medical VLMs (zero-shot), and fine-tuned VLMs (SFT vs. GRPO). Results demonstrate that our compact GRPO-tuned 2B model surpasses even 36× larger zero-shot models and achieves stronger cross-task generalization compared to SFT. Meanwhile, in each column, the best and second-best performances are marked in red and blue.
Methods Types Anatomy Disease Lesion Modality Other Overall
Identification Diagnosis Grading Recognition Attributes
Zero-Shot VLM
[…]
$\text{InstructBLIP}^{\dagger}$ [10] 44.35 32.29 59.25 75.27 23.72 40.40
[…]
Qwen2-VL-2B [2] 30.70 36.53 43.58 59.90 42.19 42.58
Qwen2-VL-7B [2] 42.57 48.83 52.06 84.74 59.66 57.57
Qwen2-VL-72B [32] 56.41 65.71 62.15 98.11 80.53 72.58
Zero-Shot Medical VLM
[…]
$\text{MedVInT}^{\dagger}$ [34] 40.26 35.78 12.77 68.10 30.30 40.04
[…]
fine-tuned VLM
Qwen2-VL-2B (SFT) 53.97 51.62 60.71 86.77 63.91 63.39
Ours-2B (GRPO) 62.88 66.08 65.87 98.24 80.14 74.64
5 Conclusion
In this work, we introduce Med-R1, a reinforcement learning (RL)-enhanced vision-language model (VLM) designed to improve medical reasoning across diverse imaging modalities. By leveraging Group Relative Policy Optimization (GRPO), Med-R1 effectively overcomes the limitations of supervised fine-tuning (SFT), significantly enhancing both cross-modality and cross-task generalization. Comprehensive evaluations across eight medical imaging modalities and five clinical reasoning tasks demonstrate that Med-R1 consistently outperforms baseline models, including zero-shot general-purpose VLMs, specialized medical VLMs, and even larger-scale models.
Despite its compact 2B parameter size, Med-R1 achieves competitive performance against 36×36$\times$36 × larger models, demonstrating that scalability does not have to come at the cost of efficiency. Its lightweight nature significantly lowers computational and deployment barriers, making it more practical for real-world clinical applications where resource constraints are critical.
While VLM-based medical reasoning remains an emerging field with ongoing challenges, Med-R1 highlights the potential of reinforcement learning in developing safer, more transparent, and clinically reliable AI-driven healthcare solutions. Future research could further enhance Med-R1 by integrating additional multi-modal clinical data, improving uncertainty quantification, and refining expert-aligned interpretability, ultimately advancing scalable and trustworthy medical reasoning systems.

### v2

@@
 Figure 1: Overview of the Evaluation Framework: Eight Medical Imaging Modalities and Five Medical Vision Question Answering Tasks.
-We evaluate Med-R1 across eight distinct medical imaging modalities—Computed Tomography (CT), Magnetic Resonance Imaging (MRI), Ultrasound, Dermoscopy, Fundus Photography, Optical Coherence Tomography (OCT), Microscopy Images, and X-ray Imaging—as well as five medical vision question answering tasks: anatomy identification, disease diagnosis, lesion grading, modality recognition, and biological attribute analysis.
+We evaluate Med-R1 across eight distinct medical imaging modalities—CT, MRI, Ultrasound, Dermoscopy, Fundus Photography, OCT, Microscopy Images, and X-ray Imaging—as well as five medical vision question answering tasks: anatomy identification, disease diagnosis, lesion grading, modality recognition, and biological attribute analysis.
 Example images and corresponding clinical questions illustrate the diversity of medical data and the challenges in developing generalizable vision-language models for automated medical reasoning.
@@
 Med-R1 exhibits superior cross-task generalization, outperforming Qwen2-VL-2B by 32.06% in question-type generalization accuracy, and even surpassing Qwen2-VL-72B in this aspect (Table 4).
-These findings emphasize the potential of RL-based training in improving both reasoning depth and task adaptability in medical vision-language models (VLMs).
-Med-R1 is a generalizable medical vision-language model (VLM) that generates clinically interpretable reasoning explanations for medical visual question answering (VQA) across eight imaging modalities (Figure 2).
-We summarized our contribution as follows:
+These findings emphasize the potential of RL-based training in improving both reasoning depth and task adaptability in medical VLMs.
+Med-R1 is a generalizable medical VLM that generates interpretable reasoning explanations for medical VQA across eight imaging modalities (Figure 2). We summarized our contribution as follows:
   1. 1.
@@
   2. 2.
-    Emergent Reasoning via Clinically Constrained RL: We adopt Group Relative Policy Optimization (GRPO), a RL strategy that eliminates the need for explicit reasoning supervision.
+    Emergent Reasoning via Clinically Constrained RL: We adopt Group Relative Policy Optimization, a RL strategy that eliminates the need for explicit reasoning supervision.
 GRPO integrates rule-based rewards (e.g., enforcing adherence to diagnostic guidelines) and group-relative comparisons, enabling models to learn clinically plausible reasoning paths from final-answer labels alone.
@@
 General VLMs and Medical VLMs.
-Vision-language models (VLMs) have significantly advanced multimodal understanding by aligning visual and textual representations through large-scale pretraining.
+VLMs have significantly advanced multimodal understanding by aligning visual and textual representations through large-scale pretraining.
 General-purpose VLMs, such as CLIP [26] and BLIP-2 [19], excel in natural image-text tasks but encounter challenges when applied to domain-specific medical reasoning.
-Recent efforts have sought to adapt these models to medical imaging through supervised fine-tuning (SFT) on domain-specific datasets.
+Recent efforts have sought to adapt these models to medical imaging through SFT on domain-specific datasets.
 For instance, LLaVA [22] introduces instruction tuning for general-purpose vision-language tasks, achieving strong performance on natural image benchmarks.
-Building on this, LLaVA-Med [18] fine-tunes LLaVA using radiology reports, demonstrating improved performance on medical visual question answering (VQA) tasks.
+Building on this, LLaVA-Med [18] fine-tunes LLaVA using radiology reports, demonstrating improved performance on medical VQA tasks.
 Similarly, Med-Flamingo [23] extends Flamingo by incorporating domain-specific pretraining for medical VQA.
 However, these SFT-based approaches often suffer from overfitting to narrow medical corpora, limiting their generalizability across diverse imaging modalities and clinical tasks.
-Our work addresses this gap by introducing reinforcement learning (RL) for generalization, enabling robust performance without modality-specific retraining.
+Our work addresses this gap by introducing RL for generalization, enabling robust performance without modality-specific retraining.
 Reinforcement Learning for Post-Training.
@@
 However, existing methods rely on complex reward models or extensive human annotation, limiting scalability in data-scarce domains like medicine.
-Our Group Relative Policy Optimization (GRPO) addresses these limitations by incorporating rule-based rewards derived from clinical guidelines and group-relative comparisons, enabling parameter-efficient adaptation to diverse medical tasks.
-Medical Reasoning and Interpretability Recent work emphasizes the importance of interpretable reasoning in medical AI, with approaches like Chain-of-Thought (CoT) prompting [33] and program-guided reasoning [6].
+The Group Relative Policy Optimization addresses these limitations by incorporating rule-based rewards derived from clinical guidelines and group-relative comparisons, enabling parameter-efficient adaptation to diverse medical tasks.
+Medical Reasoning and Interpretability Recent work emphasizes the importance of interpretable reasoning in medical AI, with approaches like CoT prompting [33] and program-guided reasoning [6].
 However, acquiring such annotations in the medical domain is particularly challenging due to the need for expert involvement [20], high annotation costs, and the complexity of medical reasoning.
-The RL-based approach enables emergent reasoning without explicit supervision, addressing this challenge.
+The RL approach enables emergent reasoning without explicit supervision, addressing this challenge.
 Figure 2: Medical VQA examples of Med-R1 The left panel demonstrates a modality recognition task, where the model correctly identifies the presence of breast tissue in a mammogram (X-ray).
@@
 Train Test CT MRI X-Ray US Der FP OCT Micro Overall
-CT 98.3 74.43 80.61 51.93 64.17 65.39 70.52 66.13 71.44
+CT 98.30 74.43 80.61 51.93 64.17 65.39 70.52 66.13 71.44
 MRI 62.33 99.29 73.75 53.18 75.04 64.57 75.59 66.31 71.26
@@
 4 Experiment & Results
-In this section, we introduce our experimental setup, implementation details, and results. We present a preliminary analysis over the results and discuss more in LABEL:sec:_discussion.
+This section introduces our experimental setup, implementation details, and results.
 4.1 Setup
 Datasets.
-We adopt the VQA data from the open-access part of the OmniMedVQA benchmark [14], which consists of a total of $82,059$ images and $88,996$ vision question answering (VQA) pairs.
+We adopt the VQA data from the open-access part of the OmniMedVQA benchmark [14], which consists of a total of $82,059$ images and $88,996$ vision question answering pairs.
 OmniMedVQA includes VQA pairs from eight imaging modalities: CT ($15,808$), MRI ($31,877$), X-Ray ($7,916$), Ultrasound ($10,991$), Dermoscopy ($6,679$), Fundus ($5,398$), OCT ($,64$ 646), and Microscopy ($5,680$).
@@
 Task setting. We evaluate our approach in two distinct generalization settings using the OmniMedVQA dataset [14]: cross-modality generalization and cross-task generalization.
-  • Cross-modality generalization: We train our model on a single modality at a time (out of the 8 available modalities) and evaluate its performance on the remaining 7 unseen modalities.
+  • Cross-modality generalization: We train our model on a single modality at a time (out of 8 available modalities) and evaluate its performance on the remaining 7 unseen modalities.
   • Cross-task generalization: We identify 5 distinct task types within the dataset and adopt the same train-test partitioning strategy as in the cross-modality setting, training on one task and evaluating on the remaining four.
 Baseline Methods & Evaluation Metric.
-Our evaluation framework integrates three classes of baselines: general VLMs including BLIP-2 [19], InstructBLIP [10], and the Qwen2-VL series (2B/7B/72B) [2, 32]; medical VLMs pretrained on domain-specific corpora (Med-Flamingo [23], MedVInT [34], LLaVA-Med [18]); and supervised fine-tuned (SFT) variants of Qwen2-VL-2B.
+Our evaluation framework integrates three classes of baselines: general VLMs including BLIP-2 [19], InstructBLIP [10], and the Qwen2-VL series (2B/7B/72B) [2, 32]; medical VLMs pretrained on domain-specific corpora (Med-Flamingo [23], MedVInT [34], LLaVA-Med [18]); and SFT variants of Qwen2-VL-2B.
 We evaluate model performance using VQA choice accuracy, the standard metric for medical VQA, where models select answers from clinically validated options.
@@
 4.2 Cross-Modality Generalization
-We comprehensively evaluate Med-R1’s adaptability across 8 medical imaging modalities, including Computed Tomography (CT), Magnetic Resonance Imaging (MRI), Ultrasound, Dermoscopy, Fundus Photography, Optical Coherence Tomography (OCT), Microscopy Images, and X-ray Imaging.
-Our experiments focus on two key aspects: (1) cross-modal generalization, where the model is trained on one modality and tested on another, and (2) comparative performance against other popular vision-language models (VLMs) and medical-specific VLMs evaluated using zero-shot and SFT.
+We comprehensively evaluate Med-R1’s adaptability across 8 medical imaging modalities, including Computed Tomography, Magnetic Resonance Imaging, Ultrasound, Dermoscopy, Fundus Photography, Optical Coherence Tomography, Microscopy Images, and X-ray Imaging.
+Our experiments focus on two key aspects: (1) cross-modal generalization, where the model is trained on one modality and tested on another, and (2) comparative performance against other popular VLMs and medical-specific VLMs evaluated using zero-shot and SFT.
 Results on generalization.
@@
 Notably, models trained on CT, MRI, and X-Ray exhibit the highest generalization capability, with overall scores of 71.44%, 71.26%, and 72.35%, respectively.
-In contrast, models trained on Fundus Photography (FP) and Microscopy (Micro) images show lower generalization, with 67.67% and 67.54% overall accuracy, indicating that certain modality-specific features (e.g., texture-based imaging in US and Micro) may not transfer as effectively to other domains.
+In contrast, models trained on Fundus Photography and Microscopy images show lower generalization, with 67.67% and 67.54% overall accuracy, indicating that certain modality-specific features (e.g., texture-based imaging in US and Micro) may not transfer as effectively to other domains.
 Importantly, the overall test accuracy of 69.91% highlights Med-R1’s ability to perform well across unseen imaging modalities, despite being trained on a single domain at a time.
-This result underscores the effectiveness of reinforcement learning (RL) in enhancing cross-modality transfer, allowing the model to maintain robust performance without requiring extensive retraining for each medical imaging modality.
+This result underscores the effectiveness of reinforcement learning in enhancing cross-modality transfer, allowing the model to maintain robust performance without requiring extensive retraining for each medical imaging modality.
 Comparison to zero-shot and SFT evaluations with other VLMs.
@@
 Modality Recognition 55.30 55.43 56.88 99.46 70.88 67.59
-Other Attributes 56.6 59.93 56.88 97.91 96.59 73.58
+Other Attributes 56.60 59.93 56.88 97.91 96.59 73.58
 Overall 62.88 66.08 65.87 98.24 80.14 74.64
@@
 5 Conclusion
-In this work, we introduce Med-R1, a reinforcement learning (RL)-enhanced vision-language model (VLM) designed to improve medical reasoning across diverse imaging modalities.
-By leveraging Group Relative Policy Optimization (GRPO), Med-R1 effectively overcomes the limitations of supervised fine-tuning (SFT), significantly enhancing both cross-modality and cross-task generalization.
+In this work, we introduce Med-R1, a RL-enhanced VLM designed to improve medical reasoning across diverse imaging modalities.
+By leveraging GRPO, Med-R1 effectively overcomes the limitations of SFT, significantly enhancing both cross-modality and cross-task generalization.
 Comprehensive evaluations across eight medical imaging modalities and five clinical reasoning tasks demonstrate that Med-R1 consistently outperforms baseline models, including zero-shot general-purpose VLMs, specialized medical VLMs, and even larger-scale models.
@@
 Its lightweight nature significantly lowers computational and deployment barriers, making it more practical for real-world clinical applications where resource constraints are critical.
-While VLM-based medical reasoning remains an emerging field with ongoing challenges, Med-R1 highlights the potential of reinforcement learning in developing safer, more transparent, and clinically reliable AI-driven healthcare solutions.
+While VLM medical reasoning remains an emerging field with ongoing challenges, Med-R1 highlights the potential of reinforcement learning in developing safer, more transparent, and clinically reliable AI-driven healthcare solutions.
 Future research could further enhance Med-R1 by integrating additional multi-modal clinical data, improving uncertainty quantification, and refining expert-aligned interpretability, ultimately advancing scalable and trustworthy medical reasoning systems.
+6 Acknowledgement
+This work is supported in part by the National Institutes of Health under award numbers R01CA272991, R01EB032680, R01DE033512, and U54CA274513.

### v3

@@
 Medical reasoning tasks, which require robust image analysis and the generation of well-justified answers, present unique challenges due to the inherent complexity of medical images.
-Transparency and trustworthiness are critical not only for gaining clinicians’ confidence but also for meeting stringent regulatory requirements.
+To be clinically useful, model outputs must be accurate, and logically consistent, as errors or inconsistencies can have critical consequences in medical applications.
 To address these challenges, we propose Med-R1, a novel framework that investigates whether reinforcement learning (RL) can enhance the generalizability and trustworthiness of VLMs in medical reasoning.
@@
 These results highlight that RL not only enhances medical reasoning capabilities but also enables parameter-efficient models to exceed the performance of significantly larger counterparts.
-Combined with its interpretable reasoning outputs, Med-R1 establishes a new potential direction for developing generalizable, trustworthy, and clinically deployable medical VLMs.
+By demonstrating strong cross-domain and cross-tasks performance, Med-R1 points toward a new direction for developing practical and generalizable medical VLMs.
 1 Introduction
 Vision-language models (VLMs) have achieved remarkable progress in reasoning over natural images (e.g., GPT-4o [16], Gemini-1.5 [33], and Qwen-VL [2]), demonstrating impressive capabilities in tasks such as visual question answering (VQA) and multimodal dialogue.
-These advancements have been driven by large-scale vision-language pretraining and supervised fine-tuning (SFT), which enable models to associate visual content with textual information.
+These advancements have been driven by large-scale vision-language pretraining and supervised fine-tuning (SFT), which enables models to associate visual content with textual information.
 However, applying VLMs to medical imaging reasoning remains a significant challenge.
-Unlike natural images, medical data demand precise interpretation, requiring VLMs to generate not only final conclusions but also step-by-step reasoning processes that align with clinical logic.
+Unlike natural images, medical data demands precise interpretation, requiring VLMs to produce not only answer outcomes but also intermediate reasoning paths that are consistent with clinical decision-making logic.
 For instance, diagnosing a lung nodule in a CT scan necessitates localizing lesions, analyzing morphological features, and integrating patient history—a reasoning chain that must be both accurate and interpretable to gain clinicians’ trust.
 Furthermore, the diversity of medical imaging modalities (e.g., CT, MRI, microscopy) and task types (e.g., diagnosis, lesion grading) poses stringent demands on model generalizability.
-These challenges raise a critical question: How can we equip VLMs with robust reasoning abilities that generalize across medical domains while maintaining clinical trustworthiness?
+These challenges raise a critical question: How can we equip VLMs with robust reasoning abilities that generalize across medical domains while maintaining medical reliability?
 In this paper, we identify that the limitations of current medical VLMs primarily stem from the inherent drawbacks of Supervised Fine-Tuning (SFT) [6, 1].
-While SFT has been widely adopted to adapt foundation models to medical imaging [23, 34, 18, 5, 21], it suffers from two fundamental issues that hinder clinical applicability.
+While SFT has been widely adopted to adapt foundation models to medical imaging [24, 36, 18, 5, 22], it suffers from two fundamental issues that hinder medical applicability.
 First, SFT inherently biases models toward memorizing task-specific shortcuts rather than learning generalizable reasoning.
@@
 As a result, existing SFT-based medical VLMs [24, 36] lack access to high-quality CoT data, leading to shallow and non-clinically grounded reasoning.
-These models frequently produce "black-box" predictions, failing to generate interpretable rationales or adapt to out-of-domain tasks (e.g., from radiology to pathology).
-This lack of transparency and robustness poses a significant challenge to clinical adoption, where explainability and reliability are paramount.
-To address these challenges, we propose Med-R1, a framework leveraging Reinforcement Learning (RL) [27] to enhance both the generalizability and trustworthiness of medical VLMs.
-Unlike SFT, RL guides models to explore diverse reasoning pathways through reward signals, encouraging them to develop intrinsic logical steps rather than relying on annotated CoT rationales.
+These models frequently produce "black-box" predictions, struggling to provide traceable reasoning or maintain performance in out-of-domain tasks.
+This lack of transparency and robustness poses a significant challenge to medical adoption, where explainability and reliability are are indispensable requirements.
+To address these challenges, we propose Med-R1, a framework leveraging Reinforcement Learning (RL) [29] to enhance both the generalizability and interpretability of medical VLMs.
+In contrast to SFT, which primarily aligns outputs with fixed supervision, RL promotes exploration of alternative reasoning strategies by leveraging reward-driven feedback, allowing models to construct internal logic without requiring explicit CoT annotations.
 This approach mitigates overfitting and shortcut learning, as demonstrated by Chu et al.
@@
 For example, rewards can be designed to prioritize reasoning steps that align with radiological decision trees, ensuring clinically plausible rationales.
-This dual mechanism not only reduces computational overhead but also enforces adherence to domain-specific constraints, making GRPO uniquely suited for medical reasoning tasks where transparency and scalability are paramount.
+This dual mechanism not only reduces computational overhead but also enforces adherence to domain-specific constraints, making GRPO uniquely suited for medical reasoning tasks where interpretability and scalability are essential for deployment.
 Figure 1: Overview of the Evaluation Framework: Eight Medical Imaging Modalities and Five Medical Vision Question Answering Tasks.
 We evaluate Med-R1 across eight distinct medical imaging modalities—CT, MRI, Ultrasound, Dermoscopy, Fundus Photography, OCT, Microscopy Images, and X-ray Imaging—as well as five medical vision question answering tasks: anatomy identification, disease diagnosis, lesion grading, modality recognition, and biological attribute analysis.
-Example images and corresponding clinical questions illustrate the diversity of medical data and the challenges in developing generalizable vision-language models for automated medical reasoning.
+Example images and corresponding medical questions illustrate the diversity of medical data and the challenges in developing generalizable vision-language models for automated medical reasoning.
 As shown in Figure 1, to rigorously validate the effectiveness of Med-R1, we conduct a comprehensive evaluation spanning eight clinically critical imaging modalities: Computed Tomography (CT), Magnetic Resonance Imaging (MRI), Ultrasound, Dermoscopy, Fundus Photography, Optical Coherence Tomography (OCT), Microscopy Images, and X-ray Imaging.
@@
 These findings emphasize the potential of RL-based training in improving both reasoning depth and task adaptability in medical VLMs.
-Med-R1 is a generalizable medical VLM that generates interpretable reasoning explanations for medical VQA across eight imaging modalities (Figure 2). We summarized our contribution as follows:
+Med-R1 is a generalizable medical VLM that generates interpretable reasoning explanations for medical VQA across eight imaging modalities (Figure 2). We summarized the contribution as follows:
   1. 1.
@@
   2. 2.
-    Emergent Reasoning via Clinically Constrained RL: We adopt Group Relative Policy Optimization, a RL strategy that eliminates the need for explicit reasoning supervision.
-GRPO integrates rule-based rewards (e.g., enforcing adherence to diagnostic guidelines) and group-relative comparisons, enabling models to learn clinically plausible reasoning paths from final-answer labels alone.
+    Reward Guided Reasoning via Constrained RL: We employ Group Relative Policy Optimization, a reinforcement learning method that removes the need for explicit reasoning annotations.
+GRPO combines rule-informed reward signals (e.g., alignment with diagnostic protocols) with group-relative optimization, allowing models to align output behavior with medically grounded reasoning from final-answer supervision alone.
   3. 3.
-    Robust Generalization with Efficiency: Med-R1 outperforms foundation model by 29.94% and SFT baselines by 15.84% in cross-modality accuracy.
-We also enhance 32.06% on foundation modal and 11.25% than SFT baselines in cross-task generalization.
-Moreover, Med-R1 surpasses other larger models in the general domain or medical domain including Qwen2-VL-72B, and MedVInT-7B, demonstrating scalability and trustworthiness for real-world deployment.
+    Robust and Efficient Generalization: Med-R1 outperforms the base foundation model by 29.94% and SFT baselines by 15.84% in cross-modality performance.
+It further improves 32.06% over the foundation model and 11.25% over SFT baselines on cross-task generalization.
+Med-R1 also exceeds much larger models in both general-domain and medical-domain settings—including Qwen2-VL-72B and MedVInT-7B—demonstrating scalable reasoning and medical transferability beyond model size.
 2 Related Works
@@
 Similarly, Med-Flamingo [24] extends Flamingo by incorporating domain-specific pretraining for medical VQA.
-However, these SFT-based approaches often suffer from overfitting to narrow medical corpora, limiting their generalizability across diverse imaging modalities and clinical tasks.
+However, these SFT-based approaches often suffer from overfitting to narrow medical corpora, limiting their generalizability across diverse imaging modalities and medical tasks.
 Our work addresses this gap by introducing RL for generalization, enabling robust performance without modality-specific retraining.
@@
 However, existing methods rely on complex reward models or extensive human annotation, limiting scalability in data-scarce domains like medicine.
-The Group Relative Policy Optimization addresses these limitations by incorporating rule-based rewards derived from clinical guidelines and group-relative comparisons, enabling parameter-efficient adaptation to diverse medical tasks.
+The Group Relative Policy Optimization addresses these limitations by incorporating rule-based rewards derived from medical guidelines and group-relative comparisons, enabling parameter-efficient adaptation to diverse medical tasks.
 Medical Reasoning and Interpretability Recent work emphasizes the importance of interpretable reasoning in medical AI, with approaches like CoT prompting [35] and program-guided reasoning [6].
@@
 The RL approach enables emergent reasoning without explicit supervision, addressing this challenge.
+Concurrent to our work, MedVLM-R1 [26] explores reinforcement learning for radiology VQA using a GRPO-based approach.
+While promising, the study is conducted on a small-scale dataset (600 training samples) and does not evaluate generalization across tasks or modalities, which makes it challenging to assess the robustness of the proposed reasoning capability and its generalizability.
+Nonetheless, this work also reflects the growing interest in applying reinforcement learning to medical vision-language models.
 Figure 2: Medical VQA examples of Med-R1 The left panel demonstrates a modality recognition task, where the model correctly identifies the presence of breast tissue in a mammogram (X-ray).
@@
 In practice, the ground truths are letter options “A, B, C, D" for multiple choice questions, and we treat all responses with correct letter options as the leading word as correct (“A…").
+3.2 No-Thinking-Med-R1
+Previous work [20] found that for the image classification task, compared with normal rule-based RL fine-tuning, excluding format accuracy and letting the model only output the final answer could lead to better performance.
+However, they only discussed this phenomenon for image classification task.
+In this paper, we further explore the thinking process of medical VQA and propose the No-Thinking-Med-R1.
 Table 1: Cross-Modality Generalization Performance of Med-R1 with RL Post-Training.
@@
 Qwen2-VL-2B (SFT) 51.74 52.83 65.57 47.65 51.91 52.26 53.99 56.58 54.07
-Ours-2B (GRPO) 66.30 71.67 77.73 57.31 72.33 71.20 71.96 70.80 69.91
+Med-R1-2B (GRPO) 66.30 71.67 77.73 57.31 72.33 71.20 71.96 70.80 69.91
 4.2 Cross-Modality Generalization
@@
 53.99%).
-Table 3: Cross-Task Generalization of Med-R1: Performance is evaluated across five clinical reasoning task types (rows: training tasks, columns: test tasks), with darker cell shading indicating stronger generalization.
+Table 3: Cross-Task Generalization of Med-R1: Performance is evaluated across five medical reasoning task types (rows: training tasks, columns: test tasks), with darker cell shading indicating stronger generalization.
 Results demonstrate that domain-specific training (e.g., Disease Diagnosis) preserves in-task expertise while maintaining adaptability to unseen tasks, particularly for modality-agnostic skills like Modality Recognition.
@@
 4.3 Cross-Task Generalization
-We also evaluate Med-R1’s generalization across 5 important clinical tasks [14]: Anatomy Identification, Disease Diagnosis, Lesion Grading, Modality Recognition, and Other Attributes.
+We also evaluate Med-R1’s generalization across 5 important medical tasks [14]: Anatomy Identification, Disease Diagnosis, Lesion Grading, Modality Recognition, and Other Attributes.
 Similar to subsection 4.2, we focus our evaluation on two aspects: cross-modality generalization and comparison against SFT and zero-shot with other VLMs.
@@
 Table 4: Performance Comparison of VLMs on Five Medical VQA Tasks: GRPO Fine-Tuning Outperforms Zero-Shot and SFT Baselines Across Diverse Reasoning Tasks.
-Performance is evaluated on five clinical reasoning types (columns) across three model categories: general-purpose VLMs (zero-shot), medical VLMs (zero-shot), and fine-tuned VLMs (SFT vs.
+Performance is evaluated on five medical reasoning types (columns) across three model categories: general-purpose VLMs (zero-shot), medical VLMs (zero-shot), and fine-tuned VLMs (SFT vs.
 GRPO).
@@
 Qwen2-VL-2B (SFT) 53.97 51.62 60.71 86.77 63.91 63.39
-Ours-2B (GRPO) 62.88 66.08 65.87 98.24 80.14 74.64
+Med-R1-2B (GRPO) 62.88 66.08 65.87 98.24 80.14 74.64
 5 Conclusion
@@
 By leveraging GRPO, Med-R1 effectively overcomes the limitations of SFT, significantly enhancing both cross-modality and cross-task generalization.
-Comprehensive evaluations across eight medical imaging modalities and five clinical reasoning tasks demonstrate that Med-R1 consistently outperforms baseline models, including zero-shot general-purpose VLMs, specialized medical VLMs, and even larger-scale models.
+Comprehensive evaluations across eight medical imaging modalities and five medical reasoning tasks demonstrate that Med-R1 consistently outperforms baseline models, including zero-shot general-purpose VLMs, specialized medical VLMs, and even larger-scale models.
 Despite its compact 2B parameter size, Med-R1 achieves competitive performance against 36×36$\times$36 × larger models, demonstrating that scalability does not have to come at the cost of efficiency.
@@
 While VLM medical reasoning remains an emerging field with ongoing challenges, Med-R1 highlights the potential of reinforcement learning in developing safer, more transparent, and clinically reliable AI-driven healthcare solutions.
-Future research could further enhance Med-R1 by integrating additional multi-modal clinical data, improving uncertainty quantification, and refining expert-aligned interpretability, ultimately advancing scalable and trustworthy medical reasoning systems.
+Future research could further enhance Med-R1 by integrating additional multi-modal medical data, improving uncertainty quantification, and refining expert-aligned interpretability, ultimately advancing scalable and trustworthy medical reasoning systems.
 6 Acknowledgement

### v4

@@
 Abstract
-Vision-language models (VLMs) have made significant progress in reasoning within natural scenes, yet their potential in medical imaging remains largely underexplored.
-Medical reasoning tasks, which require robust image analysis and the generation of well-justified answers, present unique challenges due to the inherent complexity of medical images.
-To be clinically useful, model outputs must be accurate, and logically consistent, as errors or inconsistencies can have critical consequences in medical applications.
-To address these challenges, we propose Med-R1, a novel framework that investigates whether reinforcement learning (RL) can enhance the generalizability and trustworthiness of VLMs in medical reasoning.
-Building on the recently introduced DeepSeek strategy, we adopt Group Relative Policy Optimization (GRPO) for RL, which encourages models to explore reasoning paths guided by reward signals.
-Unlike supervised fine-tuning (SFT), which often overfits to training data and struggles with generalization, RL enables models to develop more robust and diverse reasoning capabilities.
-We comprehensively evaluate Med-R1 across eight distinct medical imaging modalities: Computed Tomography (CT), Magnetic Resonance Imaging (MRI), Ultrasound, Dermoscopy, Fundus Photography, Optical Coherence Tomography (OCT), Microscopy Images, and X-ray Imaging.
-Compared to the base model, Qwen2-VL-2B, Med-R1 achieves a 29.94% improvement in average accuracy across these modalities and even outperforms Qwen2-VL-72B—a model with 36 times more parameters.
-To assess model’s generalization abilities, we further test on five different question types: modality recognition, anatomy identification, disease diagnosis, lesion grading, and biological attribute analysis.
-Med-R1 also demonstrates superior cross-task generalization, outperforming Qwen2-VL-2B by 32.06% and Qwen2-VL-72B in question-type generalization accuracy.
-These results highlight that RL not only enhances medical reasoning capabilities but also enables parameter-efficient models to exceed the performance of significantly larger counterparts.
-By demonstrating strong cross-domain and cross-tasks performance, Med-R1 points toward a new direction for developing practical and generalizable medical VLMs.
+Vision-language models (VLMs) have achieved impressive progress in natural image reasoning, yet their potential in medical imaging remains underexplored.
+Medical vision-language tasks demand precise understanding and clinically coherent answers, which are difficult to achieve due to the complexity of medical data and the scarcity of high-quality expert annotations.
+These challenges limit the effectiveness of conventional supervised fine-tuning (SFT) and Chain-of-Thought (CoT) strategies that work well in general domains.
+To address these challenges, we propose Med-R1, a reinforcement learning (RL)-enhanced vision-language model designed to improve generalization and reliability in medical reasoning.
+Built on the DeepSeek strategy, Med-R1 adopts Group Relative Policy Optimization (GRPO) to encourage reward-guided learning beyond static annotations.
+We comprehensively evaluate Med-R1 across eight distinct medical imaging modalities.
+Med-R1 achieves a 29.94% improvement in average accuracy over its base model Qwen2-VL-2B, and even outperforms Qwen2-VL-72B—a model with 36$\times$ more parameters.
+To assess cross-task generalization, we further evaluate Med-R1 on five question types.
+Med-R1 outperforms Qwen2-VL-2B by 32.06% in question-type generalization, also surpassing Qwen2-VL-72B.
+We further explore the thinking process in Med-R1, a crucial component for the success of Deepseek-R1.
+Our results show that omitting intermediate rationales (No-Thinking-Med-R1) not only improves in-domain and cross-domain generalization with less training, but also challenges the assumption that more reasoning always helps.
+These findings suggest that in medical VQA, it is not reasoning itself, but its quality and domain alignment, that determine effectiveness.
+Together, these results highlight that RL improves medical reasoning and generalization, enabling efficient and reliable VLMs for real-world deployment.
+{IEEEkeywords}
+Reinforcement Learning, Vision Language Models, Multimodal LLM, Post-Training, Medical Reasoning.
 1 Introduction
-Vision-language models (VLMs) have achieved remarkable progress in reasoning over natural images (e.g., GPT-4o [16], Gemini-1.5 [33], and Qwen-VL [2]), demonstrating impressive capabilities in tasks such as visual question answering (VQA) and multimodal dialogue.
-These advancements have been driven by large-scale vision-language pretraining and supervised fine-tuning (SFT), which enables models to associate visual content with textual information.
-However, applying VLMs to medical imaging reasoning remains a significant challenge.
-Unlike natural images, medical data demands precise interpretation, requiring VLMs to produce not only answer outcomes but also intermediate reasoning paths that are consistent with clinical decision-making logic.
-For instance, diagnosing a lung nodule in a CT scan necessitates localizing lesions, analyzing morphological features, and integrating patient history—a reasoning chain that must be both accurate and interpretable to gain clinicians’ trust.
-Furthermore, the diversity of medical imaging modalities (e.g., CT, MRI, microscopy) and task types (e.g., diagnosis, lesion grading) poses stringent demands on model generalizability.
-These challenges raise a critical question: How can we equip VLMs with robust reasoning abilities that generalize across medical domains while maintaining medical reliability?
+[…]
+Vision-language models (VLMs) have achieved strong performance on natural image understanding tasks such as visual question answering (VQA) and multimodal dialogue [1, 2, 3], enabled by large-scale pretraining and supervised fine-tuning (SFT).
+However, applying VLMs to medical imaging remains challenging due to the need for clinically sound interpretations and decision-making processes.
+Medical tasks often involve multi-step analysis—e.g., diagnosing a lung nodule may require integrating lesion localization, morphology, and context.
+In addition, the diversity of imaging modalities (e.g., CT, MRI) and task types (e.g., diagnosis, grading) imposes demands on generalizability.
+This raises a key question: how can we enable VLMs to perform well across medical domains while ensuring reliable and context-aware behavior?
 In this paper, we identify that the limitations of current medical VLMs primarily stem from the inherent drawbacks of Supervised Fine-Tuning (SFT) [4, 5].
@@
 However, curating such CoT datasets is prohibitively expensive, as it demands meticulous annotation by experienced medical professionals to ensure diagnostic validity and clinical coherence.
-As a result, existing SFT-based medical VLMs [24, 36] lack access to high-quality CoT data, leading to shallow and non-clinically grounded reasoning.
+As a result, existing SFT-based medical VLMs [6, 7] lack access to high-quality CoT data, leading to shallow reasoning with limited clinical rigor.
 These models frequently produce "black-box" predictions, struggling to provide traceable reasoning or maintain performance in out-of-domain tasks.
-This lack of transparency and robustness poses a significant challenge to medical adoption, where explainability and reliability are are indispensable requirements.
-To address these challenges, we propose Med-R1, a framework leveraging Reinforcement Learning (RL) [29] to enhance both the generalizability and interpretability of medical VLMs.
-In contrast to SFT, which primarily aligns outputs with fixed supervision, RL promotes exploration of alternative reasoning strategies by leveraging reward-driven feedback, allowing models to construct internal logic without requiring explicit CoT annotations.
-This approach mitigates overfitting and shortcut learning, as demonstrated by Chu et al.
-[9], who showed that RL-trained vision-language models significantly outperform SFT in cross-domain generalization.
-Building on recent advancements in RL optimization, we adopt Group Relative Policy Optimization (GRPO) [30], a novel strategy that eliminates the need for complex value models required by traditional methods like Proximal Policy Optimization (PPO) [29].
-GRPO incorporates rule-based rewards derived from medical guidelines (e.g., diagnostic criteria for tumors) and group-relative comparisons to stabilize policy updates.
-For example, rewards can be designed to prioritize reasoning steps that align with radiological decision trees, ensuring clinically plausible rationales.
-This dual mechanism not only reduces computational overhead but also enforces adherence to domain-specific constraints, making GRPO uniquely suited for medical reasoning tasks where interpretability and scalability are essential for deployment.
+This lack of transparency and robustness poses a significant challenge to medical adoption, where explainability and reliability are indispensable requirements.
+To address these challenges, we propose Med-R1, a reinforcement learning (RL)-based framework for enhancing the generalizability and interpretability of medical VLMs.
+Unlike SFT, which aligns outputs to fixed supervision and often leads to shortcut learning, RL encourages exploration of diverse reasoning strategies through reward signals—without requiring explicit CoT annotations [11].
+We adopt Group Relative Policy Optimization (GRPO) [12], a lightweight alternative to PPO [13] that stabilizes training via rule-based rewards and group-relative comparisons.
+These mechanisms reduce computational overhead while encouraging clinically grounded reasoning, making GRPO well-suited for medical tasks where scalability and reliability are essential.
 Figure 1: Overview of the Evaluation Framework: Eight Medical Imaging Modalities and Five Medical Vision Question Answering Tasks.
 We evaluate Med-R1 across eight distinct medical imaging modalities—CT, MRI, Ultrasound, Dermoscopy, Fundus Photography, OCT, Microscopy Images, and X-ray Imaging—as well as five medical vision question answering tasks: anatomy identification, disease diagnosis, lesion grading, modality recognition, and biological attribute analysis.
-Example images and corresponding medical questions illustrate the diversity of medical data and the challenges in developing generalizable vision-language models for automated medical reasoning.
-As shown in Figure 1, to rigorously validate the effectiveness of Med-R1, we conduct a comprehensive evaluation spanning eight clinically critical imaging modalities: Computed Tomography (CT), Magnetic Resonance Imaging (MRI), Ultrasound, Dermoscopy, Fundus Photography, Optical Coherence Tomography (OCT), Microscopy Images, and X-ray Imaging.
-These modalities were selected to represent the spectrum of medical imaging challenges, from macroscopic anatomical visualization (CT/MRI) to cellular-level analysis (Microscopy) and functional assessments (OCT).
-Across these domains, Med-R1 (2B parameters) achieves an average accuracy of 69.91%, representing a 29.94% absolute improvement over its foundation model Qwen2-VL-2B.
-Notably (Table 2), Med-R1 even surpasses Qwen2-VL-72B—a 72B-parameter model with 36× larger scale, demonstrating that parameter efficiency need not compromise performance when leveraging RL-driven reasoning.
-Beyond evaluating performance across imaging modalities, we also examine Med-R1’s generalizability across different medical question types.
-We assess the model’s capability in five key question categories: modality recognition, anatomy identification, disease diagnosis, lesion grading, and biological attribute analysis.
-Med-R1 exhibits superior cross-task generalization, outperforming Qwen2-VL-2B by 32.06% in question-type generalization accuracy, and even surpassing Qwen2-VL-72B in this aspect (Table 4).
-These findings emphasize the potential of RL-based training in improving both reasoning depth and task adaptability in medical VLMs.
-Med-R1 is a generalizable medical VLM that generates interpretable reasoning explanations for medical VQA across eight imaging modalities (Figure 2). We summarized the contribution as follows:
+Example images and corresponding clinical questions illustrate the diversity of medical data and the challenges in developing generalizable vision-language models for automated medical reasoning.
+As shown in Figure 1, we evaluate Med-R1 across eight diverse medical imaging modalities, including CT, MRI, Ultrasound, Dermoscopy, Fundus Photography, OCT, Microscopy, and X-ray.
+These modalities cover a wide range of clinical imaging—from macroscopic anatomy to cellular-level and functional assessments.
+Med-R1 (2B parameters) achieves 69.91% average accuracy, a 29.94% gain over its base model Qwen2-VL-2B, and even outperforms the 72B-parameter Qwen2-VL-72B (Table 3), highlighting the benefit of RL-driven adaptation.
+We further assess generalizability across five question types: modality recognition, anatomy identification, disease diagnosis, lesion grading, and biological attribute analysis.
+Med-R1 improves question-type generalization accuracy by 32.06% over Qwen2-VL-2B and surpasses Qwen2-VL-72B in this setting as well (Table 6).
+These demonstrate that RL enhances both parameter efficiency and task generalizability in medical VLMs.
+Finally, inspired by recent findings on reasoning behavior in rule-based RL for classification tasks [14], we further investigate the role of intermediate rationales in Med-R1.
+A widely held assumption in VLM development is that generating step-by-step reasoning (“Think”) enhances generalization and performance.
+However, our empirical analysis reveals that this may not hold in domain-specific settings like medical VQA.
+We compare two RL post-training strategies: one that guides the model to produce explicit rationales before answering, and another that directly outputs the final answer without any intermediate reasoning steps.
+Surprisingly, we observe that the latter (No-Thinking) strategy often yields better generalization across both imaging modalities and question types.
+We hypothesize that in the absence of high-quality CoT supervision, RL-trained models fall back on their pre-trained reasoning patterns, which are typically tuned for general-domain tasks.
+Under large domain shifts—such as differences in image characteristics, task goals, and clinical language—this can result in hallucinated or misaligned rationales that impair performance.
+These findings suggest that it is not the presence of reasoning per se, but rather its quality and domain alignment, that determine its utility in medical vision-language reasoning.
+Med-R1 is a generalizable medical VLM that generates interpretable reasoning explanations for medical VQA across eight imaging modalities (Figure 2). We summarized our contribution as follows:
   1. 1.
-    Generalizable Medical VLM with Multi-Modal Explicit Reasoning: We propose Med-R1, the first vision-language model supporting eight medical imaging modalities (CT, MRI, Ultrasound, etc.), capable of generating step-by-step reasoning explanations without task-specific retraining.
-Unlike prior SFT-based models limited to single modalities and black-box answers, Med-R1 achieves cross-domain generalization through RL-driven exploration.
+    Multi-modality multi-task reasoning medical VLM.
+We propose Med-R1, the first reasoning medical VLM capable of supporting eight medical imaging modalities (CT, MRI, Ultrasound, etc.) across five unique tasks.
+We demonstrate that RL-based fine-tuning effectively promotes modality-specific as well as cross-modality reasoning in the medical domain without the need for token-level supervision as in SFT.
+The proposed Med-R1 is capable of generating step-by-step, accurate, and clinically plausible explanations.
   2. 2.
-    Reward Guided Reasoning via Constrained RL: We employ Group Relative Policy Optimization, a reinforcement learning method that removes the need for explicit reasoning annotations.
-GRPO combines rule-informed reward signals (e.g., alignment with diagnostic protocols) with group-relative optimization, allowing models to align output behavior with medically grounded reasoning from final-answer supervision alone.
+    Robust Generalization with Efficiency.
+We show that alongside the solid modality and task-specific performance, Med-R1 exhibits strong generalization.
+Med-R1 outperforms the base model by 29.94% and SFT baselines by 15.84% in average generalization accuracy across modalities.
+In cross-task settings, it outperforms the base model and SFT baseline by 32.06% and 11.25% respectively.
+Moreover, Med-R1 surpasses other larger generic or medical-specific models including Qwen2-VL-72B, and MedVInT-7B, warranting its efficiency and reliability for real-world deployment.
   3. 3.
-    Robust and Efficient Generalization: Med-R1 outperforms the base foundation model by 29.94% and SFT baselines by 15.84% in cross-modality performance.
-It further improves 32.06% over the foundation model and 11.25% over SFT baselines on cross-task generalization.
-Med-R1 also exceeds much larger models in both general-domain and medical-domain settings—including Qwen2-VL-72B and MedVInT-7B—demonstrating scalable reasoning and medical transferability beyond model size.
+    Rethinking the “More Thinking is Better” Assumption: Our results challenge the common belief that generating explicit thinking processes improves generalization.
+In medical VQA, we find that RL without reasoning steps yields more accurate predictions, while free-form reasoning—learned from general-domain data—can lead to hallucinations under domain shift.
+This highlights that in specialized domains—where image types, task goals, and language style differ significantly from general VLM training—the quality of reasoning, rather than its quantity, is critical for generalization.
 2 Related Works
 General VLMs and Medical VLMs.
-VLMs have significantly advanced multimodal understanding by aligning visual and textual representations through large-scale pretraining.
-General-purpose VLMs, such as CLIP [28] and BLIP-2 [19], excel in natural image-text tasks but encounter challenges when applied to domain-specific medical reasoning.
-Recent efforts have sought to adapt these models to medical imaging through SFT on domain-specific datasets.
-For instance, LLaVA [23] introduces instruction tuning for general-purpose vision-language tasks, achieving strong performance on natural image benchmarks.
-Building on this, LLaVA-Med [18] fine-tunes LLaVA using radiology reports, demonstrating improved performance on medical VQA tasks.
-Similarly, Med-Flamingo [24] extends Flamingo by incorporating domain-specific pretraining for medical VQA.
-However, these SFT-based approaches often suffer from overfitting to narrow medical corpora, limiting their generalizability across diverse imaging modalities and medical tasks.
-Our work addresses this gap by introducing RL for generalization, enabling robust performance without modality-specific retraining.
+General-purpose VLMs such as CLIP [15] and BLIP-2 [16] have advanced natural image-text understanding via large-scale pretraining, but struggle with domain-specific tasks like medical reasoning.
+Recent efforts adapt VLMs through supervised fine-tuning (SFT) on medical datasets, as seen in LLaVA-Med [8] and Med-Flamingo [6].
+While effective in-domain, these models often overfit to narrow corpora and lack generalization across modalities or task types.
+Our work addresses this limitation by introducing RL for scalable, modality-agnostic adaptation.
 Reinforcement Learning for Post-Training.
-Reinforcement learning [32, 8] has emerged as a powerful tool for aligning large language models (LLMs) with human preferences through reward signals.
-Approaches like RLHF [25] and RLAIF [4] leverage human or AI-generated feedback to refine model behavior.
-In the vision-language domain, RL has been applied to improve VQA accuracy [9] and reduce hallucination [30].
-However, existing methods rely on complex reward models or extensive human annotation, limiting scalability in data-scarce domains like medicine.
-The Group Relative Policy Optimization addresses these limitations by incorporating rule-based rewards derived from medical guidelines and group-relative comparisons, enabling parameter-efficient adaptation to diverse medical tasks.
-Medical Reasoning and Interpretability Recent work emphasizes the importance of interpretable reasoning in medical AI, with approaches like CoT prompting [35] and program-guided reasoning [6].
-However, acquiring such annotations in the medical domain is particularly challenging due to the need for expert involvement [21], high annotation costs, and the complexity of medical reasoning.
-The RL approach enables emergent reasoning without explicit supervision, addressing this challenge.
-Concurrent to our work, MedVLM-R1 [26] explores reinforcement learning for radiology VQA using a GRPO-based approach.
-While promising, the study is conducted on a small-scale dataset (600 training samples) and does not evaluate generalization across tasks or modalities, which makes it challenging to assess the robustness of the proposed reasoning capability and its generalizability.
-Nonetheless, this work also reflects the growing interest in applying reinforcement learning to medical vision-language models.
-Figure 2: Medical VQA examples of Med-R1 The left panel demonstrates a modality recognition task, where the model correctly identifies the presence of breast tissue in a mammogram (X-ray).
+RL has shown promise for aligning language models with desired behavior via reward feedback [17, 18].
+In vision-language tasks, RL improves VQA accuracy [11] and reduces hallucination [12], but often relies on complex reward models or costly human supervision.
+GRPO [12] offers a scalable alternative by using rule-based rewards and group-relative comparisons.
+We extend this approach to medical VQA, enabling efficient adaptation without modality-specific supervision.
+Medical Reasoning and Interpretability Interpretable reasoning is critical in medical AI, with recent work exploring CoT prompting [19] and program-guided logic [4].
+However, such methods rely on costly expert annotations [20], limiting scalability in clinical domains.
+Reinforcement learning offers an alternative by enabling emergent reasoning without explicit supervision.
+Concurrently, MedVLM-R1 [21] also applies GRPO-based RL to radiology VQA, which is conducted on a relatively small dataset without generalization analysis, making its reasoning robustness challenging to assess.
+Nonetheless, it highlights the growing interest in RL for medical VLMs.
+Figure 2: Medical VQA examples of Med-R1 The left panel shows a modality recognition task, where the model correctly identifies the presence of breast tissue in a mammogram (X-ray).
 The right panel illustrates a disease diagnosis task, where Med-R1 accurately detects urolithiasis in an abdominal CT scan.
-In both cases, the model provides interpretable reasoning (“Think”), explaining its decision-making process before selecting the final answer.
+In both cases, the model provides reasoning (“Think”), explaining its decision-making process before selecting the final answer.
 3 Method
-We adopt RL to encourage multimodal reasoning and generalization in medical domains.
+We adopt rule-based RL to encourage multimodal reasoning and generalization in medical domains.
 Recent studies have demonstrated that RL can incentivize emergent logical reasoning and generalization in multimodal tasks such as mathematical reasoning [22] and visual navigation [11].
@@
 3.1 Group Relative Policy Optimization (GRPO)
-Overview.
-RL-based algorithms such as PPO [29] and GRPO [30] belong to a family of fine-tuning and alignment strategies explicitly designed to enhance models’ reasoning capacities.
+Overview: RL-based algorithms such as PPO [13] and GRPO [12] belong to a family of fine-tuning and alignment strategies explicitly designed to enhance models’ reasoning capacities.
 Unlike supervised fine-tuning, which directly optimizes maximum likelihood, these RL-based methods instead optimize the policy gradient using reward signals, encouraging reasoning by exploring a much larger solution space.
@@
 These optimizations make GRPO 50% more resource- and computation-efficient than PPO [12].
-Definition.
-Formally, let $P(Q)$ denote the question set used for training, where $q$ is a sampled question in the current iteration.
+Definition: Formally, let $P(Q)$ denote the question set used for training, where $q$ is a sampled question in the current iteration.
 Let $\pi_{\theta_{\text{old}}}$ and $\pi_{\theta_{\text{new}}}$ denote the old policy and current (new) policy, respectively, where $o$ is a complete response sampled from a policy.
@@
 The GRPO objective is given by:
-$\mathcal{J}_{\text{GRPO}}(\theta)$ $=\mathbb{E}_{q\sim P(Q),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O|q)}$ (1)
-$\frac{1}{G}\sum_{i=1}^{G}\Bigg{(}\min\Bigg{(}\frac{\pi_{\theta_{\text{new}}}(o_{i}|q)}{\pi_{\theta_{\text{old}}}(o_{i}|q)}A_{i},\text{clip}\Bigg{(}\frac{\pi_{\theta_{\text{new}}}(o_{i}|q)}{\pi_{\theta_{\text{old}}}(o_{i}|q)},1-\epsilon,1+\epsilon\Bigg{)}A_{i}\Bigg{)}-\beta\mathbb{D}_{KL}\Big{(}\pi_{\theta_{\text{new}}}\Big{\|}\pi_{\text{ref}}\Big{)}\Bigg{)}$
-where $\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\text{old}}}(o_{i}|q)}$ is the policy ratio, and $A_{i}$ is the estimated advantage.
-The KL divergence term [17] regularizes the policy update, ensuring that $\pi_{\theta}$ does not deviate excessively from the reference model $\pi_{\theta_{\text{ref}}}$.
+$\mathcal{J}_{\text{GRPO}}(\theta)=$ $\ \mathbb{E}_{q\sim P(Q),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O\mid q)}$
+$\quad\frac{1}{G}\sum_{i=1}^{G}\Bigg{[}\min\Bigg{(}\frac{\pi_{\theta_{\text{new}}}(o_{i}\mid q)}{\pi_{\theta_{\text{old}}}(o_{i}\mid q)}A_{i},$
+$\qquad\text{clip}\left(\frac{\pi_{\theta_{\text{new}}}(o_{i}\mid q)}{\pi_{\theta_{\text{old}}}(o_{i}\mid q)},1-\epsilon,1+\epsilon\right)A_{i}\Bigg{)}$
+$\quad\quad-\beta\,\mathbb{D}_{\text{KL}}\Big{(}\pi_{\theta_{\text{new}}}\Big{\|}\pi_{\text{ref}}\Big{)}\Bigg{]}$ (1)
+where $\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\text{old}}}(o_{i}|q)}$ is the policy ratio, and $A_{i}$ is the estimated advantage, and $\epsilon$ is the clipping threshold for policy updates The KL divergence term [25] regularizes the policy update, ensuring that $\pi_{\theta}$ does not deviate excessively from the reference model $\pi_{\theta_{\text{ref}}}$.
 Unlike PPO, which uses a critic model to estimate the advantage $A_{i}$ for a single response $o$, GRPO estimates the relative advantage by sampling a group of responses $\{o_{i}\}_{i=1}^{G}$ and normalizing their rewards within the group to compute a relative advantage [12, 26].
-Each reward is calculated based on rules without using a regard model.
-We detail the reward design below.
-Reward design.
-We follow [30] and use two types of reward: format and accuracy.
-To begin with, we prompt the model to explicitly output its thinking process in the “<think>…</think>" tag and the final answer in the “<answer>…</answer>" tag.
-The format reward is designed to check if the aforementioned tags are present in the final response.
-A reward score of 1 will be given if they exist and are correct.
-This helps the model to organize its thoughts and answer in a structured format for the ease of reading.
-The accuracy reward is a rule-based reward that checks if the actually answer matches with the ground truth.
-Similarly, a reward score of 1 is given when the results match.
-In practice, the ground truths are letter options “A, B, C, D" for multiple choice questions, and we treat all responses with correct letter options as the leading word as correct (“A…").
-3.2 No-Thinking-Med-R1
-Previous work [20] found that for the image classification task, compared with normal rule-based RL fine-tuning, excluding format accuracy and letting the model only output the final answer could lead to better performance.
-However, they only discussed this phenomenon for image classification task.
-In this paper, we further explore the thinking process of medical VQA and propose the No-Thinking-Med-R1.
+Each reward is calculated based on rules without reward models.
+We detail reward design below.
 Table 1: Cross-Modality Generalization Performance of Med-R1 with RL Post-Training.
 Accuracy (%) across eight medical imaging modalities, where rows indicate training modalities and columns test modalities.
-Darker cell shades indicate higher accuracy for corresponding training-test pairs.
-The "Overall" column/row reports average performance per training and test domain.
+Darker cell shades indicate higher accuracy for corresponding training-test pairs in each column.
+The base model is Qwen2-VL-2B.
 RL fine-tuned VLM
@@
 Overall 66.30 71.67 77.73 57.31 72.33 71.20 71.96 70.80 69.91
-  • CT - Computed Tomography; MRI - Magnetic Resonance Imaging; US - Ultrasound.
-  • Der - Dermoscopy; FP - Fundus Photography; OCT - Optical Coherence Tomography
-  • Micro - Microscopy Images; X-Ray - X-Ray Imaging
+  • CT - Computed Tomography; MRI - Magnetic Resonance Imaging; US - Ultrasound; Der - Dermoscopy; FP - Fundus Photography.
+  • OCT - Optical Coherence Tomography; Micro - Microscopy Images; X-Ray - X-Ray Imaging
+Table 2: Cross-Modality Generalization Performance of Med-R1 with RL Post-Training.
+Accuracy (%) across eight medical imaging modalities, where rows indicate training modalities and columns test modalities.
+Darker cell shades indicate higher accuracy for corresponding training-test pairs in each column.
+The base model is Qwen2-VL-2B.
+No-Think RL fine-tuned VLM
+Train Test CT MRI X-Ray US Der FP OCT Micro Overall
+CT 99.54 82.18 80.00 46.24 72.13 72.50 76.06 70.36 74.88
+MRI 79.27 99.72 74.74 42.86 79.94 73.50 84.67 69.10 75.48
+X-Ray 78.37 69.94 97.09 46.43 70.75 80.97 69.81 70.63 73.00
+US 59.82 66.53 73.37 100.00 70.60 68.03 68.63 79.28 73.28
+Der 56.34 67.94 73.87 46.14 98.00 69.67 67.45 67.57 68.37
+FP 59.55 68.62 73.93 51.59 72.43 93.99 72.64 70.36 70.39
+OCT 80.47 73.50 77.15 49.08 65.08 70.95 100.00 67.21 72.93
+Micro 64.15 66.56 76.84 53.13 68.91 70.95 73.47 98.56 71.50
+Overall 72.19 74.37 78.37 54.43 74.73 75.07 76.59 74.13 72.48
+Reward design: We follow [12] and use two types of reward: format and accuracy.
+Firstly, we prompt the model to explicitly output its thinking process in the “<think>…</think>" tag and the final answer in the “<answer>…</answer>" tag.
+The format reward is designed to check if the aforementioned tags are present in the final response.
+A reward score of 1 will be given if they exist and are correct.
+This helps the model to organize its thoughts and answer in a structured format for the ease of reading.
+The accuracy reward is a rule-based reward that checks if the actual answer matches with the ground truth.
+Similarly, a reward score of 1 is given when the results match.
+In practice, the ground truths are letter options “A, B, C, D" for multiple choice questions, and we treat all responses with correct letter options as the leading word as correct (“A…").
+3.2 No-Thinking-Med-R1
+Previous work [14] found that, in image classification tasks, removing both reasoning and format supervision during RL could sometimes improve performance.
+However, that finding is specific to structured classification tasks with simple output space, where reasoning is rarely required.
+In contrast, medical VQA involves multimodal, semantically complex inputs, and reasoning failures can arise from domain mismatch rather than verbosity alone [12, 6].
+Given the difficulty of obtaining reliable CoT annotations in medical settings, we further investigate the role of intermediate reasoning in RL post-training.
+We revise the instruction prompt as {Question}.
+Output the single-letter choice (A, B, C, D,...) in <answer>...</answer> tags., where {Question} will be replaced by each specific question.
+By doing so, we encourage the model only to output the final answer without any explicit thinking process.
+The accuracy reward is maintained, where the reward score is 1 when the extracted answer matches the ground truth labels.
+Note that when there is any text outside the <answer> tag, i.e., an explicit thinking process exists, the extracted content will be null, and therefore the accuracy reward will be 0.
+Therefore, the model will be forced to generate only the answers.
+Table 3: Performance comparison across eight medical modalities.
+Our GRPO-finetuned model outperforms zero-shot general-purpose VLMs, medical-domain VLMs, and supervised fine-tuning baselines, while maintaining scalability.
+The best and second-best performances per column are highlighted in red and blue.
+Modality abbreviations: CT – Computed Tomography; MRI – Magnetic Resonance Imaging; US – Ultrasound; Der – Dermoscopy; FP – Fundus Photography; OCT – Optical Coherence Tomography; Micro – Microscopy; X-Ray – X-ray Imaging.
+Methods Modality CT MRI X-Ray US Der FP OCT Micro Overall
+Zero-shot VLMs
+[…]
+$\text{InstructBLIP}^{\dagger}$ [27] 28.72 33.15 61.04 41.25 62.22 50.31 42.59 46.29 45.70
+$\text{LLaVA}^{\dagger}$ [28] 17.73 26.72 30.70 18.66 49.74 47.11 33.73 28.87 31.66
+[…]
+[…]
+Qwen2-VL-2B [3] 45.10 38.57 39.32 30.86 35.83 43.17 35.14 36.85 38.11
+Qwen2-VL-7B [3] 61.46 45.77 64.27 36.01 49.08 59.84 59.32 61.08 54.60
+Qwen2-VL-72B [30] 67.97 69.39 77.21 51.39 65.31 72.58 72.76 67.83 68.05
+Qwen2.5-VL-3B [31] 53.87 54.23 61.84 32.69 52.94 62.47 56.23 59.64 54.24
+Qwen2.5-VL-7B [31] 60.44 58.44 73.99 30.66 62.48 67.30 61.20 67.84 60.29
+Qwen2.5-VL-72B [31] 66.18 68.74 77.59 49.81 69.75 71.04 69.22 69.37 67.71
+Zero-shot Medical VLMs
+[…]
+$\text{RadFM}^{\dagger}$ [32] 27.56 24.06 30.95 16.57 39.21 36.89 32.80 27.97 29.50
+[…]
+$\text{MedVInT}^{\dagger}$ [7] 40.74 43.10 55.10 41.26 29.11 31.84 23.26 32.00 37.05
+Fine-tuned VLMs
+Qwen2-VL-2B (SFT) 51.74 52.83 65.57 47.65 51.91 52.26 53.99 56.58 54.07
+Qwen2.5-VL-3B (SFT) 56.06 60.81 69.23 41.77 60.11 69.19 63.95 65.66 60.85
+Qwen2-VL-2B (GRPO + Think) 66.30 71.67 77.73 57.31 72.33 71.20 71.96 70.80 69.91
+Qwen2-VL-2B (GRPO + Nothink) 72.19 74.37 78.37 54.43 74.73 75.07 76.59 74.13 72.49
+Qwen2.5-VL-3B (GRPO + Nothink) 69.89 72.91 84.52 43.91 73.62 80.10 84.18 71.44 72.57
 4 Experiment & Results
-This section introduces our experimental setup, implementation details, and results.
 4.1 Setup
@@
 The GRPO policy generates four candidate rationales per sample, with a sampling temperature of $\tau=0.7$.
-Each training configuration is run for one epoch.
+Each training is run for one epoch.
 Task setting. We evaluate our approach in two distinct generalization settings using the OmniMedVQA dataset [33]: cross-modality generalization and cross-task generalization.
-  • Cross-modality generalization: We train our model on a single modality at a time (out of 8 available modalities) and evaluate its performance on the remaining 7 unseen modalities.
-  • Cross-task generalization: We identify 5 distinct task types within the dataset and adopt the same train-test partitioning strategy as in the cross-modality setting, training on one task and evaluating on the remaining four.
+  • Cross-modality generalization: We train our model on a single modality at a time (out of 8 available modalities) and evaluate its performance on the other seven modalities.
+  • Cross-task generalization: We identify 5 distinct task types within the dataset and adopt the same train-test partitioning strategy as in the cross-modality setting, training on one task and evaluating on the other four.
 Baseline Methods & Evaluation Metric.
@@
 where $N$ is the total number of test cases, $\hat{y}_{i}$ is the model’s predicted answer index, $y_{i}$ is the ground truth index, and $\mathbb{I}$ is the indicator function, which returns 1 if the predicted answer matches the ground truth and 0 otherwise.
-Table 2: Performance Comparison of VLMs on Eight Modality: Performance comparison across eight medical imaging modalities demonstrates that our GRPO-finetuned model surpasses zero-shot general-purpose VLMs, domain-specific medical VLMs, and conventional supervised fine-tuning approaches, achieving superior accuracy while maintaining scalability for clinical deployment.
-Meanwhile, in each column, the best and second-best performances are marked in red and blue.
-Naming details are as follows: CT - Computed Tomography; MRI - Magnetic Resonance Imaging; US - Ultrasound; Der - Dermoscopy; FP - Fundus Photography; OCT - Optical Coherence Tomography; Micro - Microscopy Images; X-Ray - X-Ray Imaging.
-Methods Modality CT MRI X-Ray US Der FP OCT Micro Overall
-Zero-shot VLM
-[…]
-$\text{InstructBLIP}^{\dagger}$ [10] 28.72 33.15 61.04 41.25 62.22 50.31 42.59 46.29 45.70
-[…]
-Qwen2-VL-2B [2] 45.10 38.57 39.32 30.86 35.83 43.17 35.14 36.85 38.11
-Qwen2-VL-7B [2] 61.46 45.77 64.27 36.01 49.08 59.84 59.32 61.08 54.60
-Qwen2-VL-72B [34] 67.97 69.39 77.21 51.39 65.31 72.58 72.76 67.83 68.05
-Zero-shot Medical VLM
-[…]
-$\text{MedVInT}^{\dagger}$ [36] 40.74 43.10 55.10 41.26 29.11 31.84 23.26 32.00 37.05
-[…]
-Fine-tuned VLM
-Qwen2-VL-2B (SFT) 51.74 52.83 65.57 47.65 51.91 52.26 53.99 56.58 54.07
-Med-R1-2B (GRPO) 66.30 71.67 77.73 57.31 72.33 71.20 71.96 70.80 69.91
 4.2 Cross-Modality Generalization
-We comprehensively evaluate Med-R1’s adaptability across 8 medical imaging modalities, including Computed Tomography, Magnetic Resonance Imaging, Ultrasound, Dermoscopy, Fundus Photography, Optical Coherence Tomography, Microscopy Images, and X-ray Imaging.
+We comprehensively evaluate Med-R1’s adaptability across eight medical imaging modalities, including Computed Tomography, Magnetic Resonance Imaging, Ultrasound, Dermoscopy, Fundus Photography, Optical Coherence Tomography, Microscopy Images, and X-ray Imaging.
 Our experiments focus on two key aspects: (1) cross-modal generalization, where the model is trained on one modality and tested on another, and (2) comparative performance against other popular VLMs and medical-specific VLMs evaluated using zero-shot and SFT.
+Table 4: Cross-Task Generalization of Med-R1: Performance is evaluated across five clinical reasoning task types (rows: training tasks, columns: test tasks), with darker cell shading indicating stronger generalization.
+Results demonstrate that domain-specific training (e.g., Disease Diagnosis) preserves in-task expertise while maintaining adaptability to unseen tasks, particularly for modality-agnostic skills like Modality Recognition.
+The base model is Qwen2-VL-2B.
+Train Test Anatomy Disease Lesion Modality Other Overall
+Identification Diagnosis Grading Recognition Attributes
+Anatomy Identification 96.06 59.07 55.51 98.62 74.86 76.83
+Disease Diagnosis 54.16 98.25 73.85 97.87 84.09 81.64
+Lesion Grading 52.30 57.72 86.24 97.32 74.29 73.57
+Modality Recognition 55.30 55.43 56.88 99.46 70.88 67.59
+Other Attributes 56.60 59.93 56.88 97.91 96.59 73.58
+Overall 62.88 66.08 65.87 98.24 80.14 74.64
+Table 5: No-Think Cross-Task Generalization of Med-R1: Performance is evaluated across five clinical reasoning task types (rows: training tasks, columns: test tasks), with darker cell shading indicating stronger generalization.
+The base model is Qwen2-VL-2B.
+Train Test Anatomy Disease Lesion Modality Other Overall
+Identification Diagnosis Grading Recognition Attributes
+Anatomy Identification 96.33 59.29 64.22 99.04 76.56 79.09
+Disease Diagnosis 55.00 98.30 63.76 98.08 85.23 80.07
+Lesion Grading 54.22 58.04 96.33 97.37 72.30 75.65
+Modality Recognition 56.84 57.95 55.05 99.79 73.15 68.56
+Other Attributes 56.29 58.00 52.29 97.90 99.29 72.75
+Overall 63.74 66.32 66.33 98.44 81.31 75.22
 Results on generalization.
 To evaluate Med-R1’s cross-modality generalization, we measure its accuracy across eight distinct medical imaging modalities (Table 1).
-The Overall row and column summarize the model’s average performance across training and test domains, providing insights into its generalization ability.
+Overall row and column summarize the model’s average performance across training and test domains, providing insights into its generalization ability.
 Med-R1 achieves a strong overall accuracy of 69.91%, demonstrating its ability to generalize across diverse medical imaging modalities.
@@
 This result underscores the effectiveness of reinforcement learning in enhancing cross-modality transfer, allowing the model to maintain robust performance without requiring extensive retraining for each medical imaging modality.
+Table 6: Performance Comparison of VLMs on Five Medical VQA Tasks: GRPO Fine-Tuning Outperforms Zero-Shot and SFT Baselines Across Diverse Reasoning Tasks.
+Performance is evaluated on five clinical reasoning types (columns) across three model categories: general-purpose VLMs (zero-shot), medical VLMs (zero-shot), and fine-tuned VLMs.
+The best and second-best performances are marked in red and blue.
+Methods Types Anatomy Disease Lesion Modality Other Overall
+Identification Diagnosis Grading Recognition Attributes
+Zero-shot VLMs
+[…]
+$\text{InstructBLIP}^{\dagger}$ [27] 44.35 32.29 59.25 75.27 23.72 40.40
+$\text{LLaVA}^{\dagger}$ [28] 25.86 29.10 43.95 21.36 31.90 27.96
+[…]
+[…]
+Qwen2-VL-2B [3] 30.70 36.53 43.58 59.90 42.19 42.58
+Qwen2-VL-7B [3] 42.57 48.83 52.06 84.74 59.66 57.57
+Qwen2-VL-72B [30] 56.41 65.71 62.15 98.11 80.53 72.58
+Qwen2.5-VL-3B [31] 35.23 50.45 52.79 85.23 54.77 55.69
+Qwen2.5-VL-7B [31] 41.00 61.32 54.13 97.78 70.45 64.94
+Qwen2.5-VL-72B [31] 57.22 62.55 60.32 98.20 77.41 71.14
+Zero-shot Medical VLMs
+[…]
+$\text{RadFM}^{\dagger}$ [32] 13.31 21.69 30.35 26.64 43.85 26.99
+[…]
+$\text{MedVInT}^{\dagger}$ [7] 40.26 35.78 12.77 68.10 30.30 40.04
+Fine-tuned VLMs
+Qwen2-VL-2B (SFT) 53.97 51.62 60.71 86.77 63.91 63.39
+Qwen2.5-VL-3B (SFT) 54.91 57.75 64.04 84.95 71.56 66.64
+Qwen2-VL-2B (GRPO + Think) 62.88 66.08 65.87 98.24 80.14 74.64
+Qwen2-VL-2B (GRPO + Nothink) 63.74 66.32 66.33 98.44 81.31 75.22
+Qwen2.5-VL-3B (GRPO + Nothink) 55.47 68.72 80.00 96.87 85.51 77.31
 Comparison to zero-shot and SFT evaluations with other VLMs.
@@
 53.99%).
-Table 3: Cross-Task Generalization of Med-R1: Performance is evaluated across five medical reasoning task types (rows: training tasks, columns: test tasks), with darker cell shading indicating stronger generalization.
-Results demonstrate that domain-specific training (e.g., Disease Diagnosis) preserves in-task expertise while maintaining adaptability to unseen tasks, particularly for modality-agnostic skills like Modality Recognition.
-Train Test Anatomy Disease Lesion Modality Other Overall
-Identification Diagnosis Grading Recognition Attributes
-Anatomy Identification 96.06 59.07 55.51 98.62 74.86 76.83
-Disease Diagnosis 54.16 98.25 73.85 97.87 84.09 81.64
-Lesion Grading 52.30 57.72 86.24 97.32 74.29 73.57
-Modality Recognition 55.30 55.43 56.88 99.46 70.88 67.59
-Other Attributes 56.60 59.93 56.88 97.91 96.59 73.58
-Overall 62.88 66.08 65.87 98.24 80.14 74.64
 4.3 Cross-Task Generalization
-We also evaluate Med-R1’s generalization across 5 important medical tasks [14]: Anatomy Identification, Disease Diagnosis, Lesion Grading, Modality Recognition, and Other Attributes.
+We also evaluate Med-R1’s generalization across five important clinical tasks [33]: Anatomy Identification, Disease Diagnosis, Lesion Grading, Modality Recognition, and Other Attributes.
 Similar to subsection 4.2, we focus our evaluation on two aspects: cross-modality generalization and comparison against SFT and zero-shot with other VLMs.
+Figure 3: Cross-modality accuracy difference between No-Thinking-Med-R1 and Med-R1.
+Each cell shows the performance gap (%) when trained on one modality (y-axis) and evaluated on others (x-axis).
+Red indicates improvement; blue indicates degradation.
+Figure 4: Cross-task accuracy difference between No-Thinking-Med-R1 and Med-R1.
+Each cell shows the performance gap (%) when trained on one task (y-axis) and tested on another (x-axis).
+Red indicates improvement; blue indicates degradation.
 Results on generalization.
@@
 In contrast, the average generalization with the identical base model trained with SFT is merely 63.39%, 11.25% below Med-R1, further demonstrating the strong generalization capability of Med-R1.
-Table 4: Performance Comparison of VLMs on Five Medical VQA Tasks: GRPO Fine-Tuning Outperforms Zero-Shot and SFT Baselines Across Diverse Reasoning Tasks.
-Performance is evaluated on five medical reasoning types (columns) across three model categories: general-purpose VLMs (zero-shot), medical VLMs (zero-shot), and fine-tuned VLMs (SFT vs.
-GRPO).
-Results demonstrate that our compact GRPO-tuned 2B model surpasses even 36× larger zero-shot models and achieves stronger cross-task generalization compared to SFT.
-Meanwhile, in each column, the best and second-best performances are marked in red and blue.
-Methods Types Anatomy Disease Lesion Modality Other Overall
-Identification Diagnosis Grading Recognition Attributes
-Zero-shot VLM
-[…]
-$\text{InstructBLIP}^{\dagger}$ [10] 44.35 32.29 59.25 75.27 23.72 40.40
-[…]
-Qwen2-VL-2B [2] 30.70 36.53 43.58 59.90 42.19 42.58
-Qwen2-VL-7B [2] 42.57 48.83 52.06 84.74 59.66 57.57
-Qwen2-VL-72B [34] 56.41 65.71 62.15 98.11 80.53 72.58
-Zero-shot Medical VLM
-[…]
-$\text{MedVInT}^{\dagger}$ [36] 40.26 35.78 12.77 68.10 30.30 40.04
-[…]
-Fine-tuned VLM
-Qwen2-VL-2B (SFT) 53.97 51.62 60.71 86.77 63.91 63.39
-Med-R1-2B (GRPO) 62.88 66.08 65.87 98.24 80.14 74.64
+Figure 5: Rewards & Length.
+The top row shows the change in RL rewards during Med-R1 training across eight imaging modalities and five task types.
+We observe that the model quickly converges to stable reward values within 100–200 steps, indicating the computational efficiency of our RL post-training process.
+The bottom row shows the output completion length over training steps.
+Interestingly, the output length gradually decreases as training progresses, suggesting that the reward function discourages overly verbose or unfaithful reasoning.
+This observation challenges the conventional assumption that longer reasoning necessarily benefits vision-language performance in VQA.
+4.4 Analysis of Thinking Process
+We present the results of No-Thinking-Med-R1 in this section.
+Cross-task performance is shown in Table 5, and cross-modality performance is shown in Table 2.
+Overall, No-Thinking-Med-R1 achieves stronger in-domain performance across all settings compared to Med-R1, suggesting that removing explicit reasoning generation may lead to more effective task-specific learning.
+To assess generalization, we compare Med-R1 and No-Thinking-Med-R1 across modalities and tasks.
+As visualized in Figure 3 and Figure 4, No-Thinking-Med-R1 consistently outperforms Med-R1 in cross-modality settings.
+In cross-task scenarios, however, the results are more mixed—showing gains in some tasks and declines in others.
+These findings reinforce the practical motivation behind Med-R1.
+In clinical domains, acquiring high-quality CoT supervision is prohibitively costly and often impractical.
+While reasoning improves performance in general domains, our results show that this does not necessarily hold in medical settings, where unsupervised rationales may become unreliable under domain shift.
+Med-R1 demonstrates that even without reasoning supervision, RL can offer a robust and efficient path for adapting VLMs to medicine.
+4.5 Analysis of Rewards and Completion Length.
+The top row of Figure 5 shows the reward dynamics during GRPO-based reinforcement learning.
+Across all modalities and tasks, Med-R1 converges within 100–200 steps, as evidenced by the log-scale reward curves.
+This rapid and consistent convergence highlights the efficiency of our training pipeline and the model’s ability to quickly adapt to the reward structure.
+Moreover, the smooth reward trajectories suggest that GRPO remains stable even with limited RL updates, making it well-suited for low-resource scenarios.
+Unlike in scientific domains such as mathematics [12], where model responses become increasingly detailed as thinking deepens, we observe a different trend in Med-RL training: the average response length decreases rapidly and then fluctuates (Figure 5).
+Rather than generating elaborate rationales, the model learns to produce shorter, more direct outputs—potentially favoring conciseness over verbosity.
+This aligns with our findings in subsection 4.4, suggesting that extended reasoning may not always benefit cross-domain VQA, especially under large domain shifts.
+We hypothesize that RL without CoT supervision encourages the model to rely on its pre-trained knowledge when generating rationales.
+While this may be effective in general domains, it can lead to hallucinated or misaligned reasoning in specialized domains like medical images.
+These results challenge the common assumption that more thinking always improves performance, and highlight that the “quality”—rather than the “presence”—of thinking is a crucial factor affecting the outcome of RL post-training.
+4.6 Limitations and Future Work.
+This work marks an initial step in applying RL to medical vision-language models.
+We adopt a frame-level VQA setting for consistent evaluation across modalities, but this simplifies the real-world complexity of medical imaging.
+In practice, CT and MRI are volumetric, and ultrasound is dynamic, requiring reasoning across slices and time.
+Future directions include extending Med-R1 to support multiple inputs.
+We also plan to incorporate patient context and explore more advanced reasoning frameworks for clinical deployment.
 5 Conclusion
-In this work, we introduce Med-R1, a RL-enhanced VLM designed to improve medical reasoning across diverse imaging modalities.
-By leveraging GRPO, Med-R1 effectively overcomes the limitations of SFT, significantly enhancing both cross-modality and cross-task generalization.
-Comprehensive evaluations across eight medical imaging modalities and five medical reasoning tasks demonstrate that Med-R1 consistently outperforms baseline models, including zero-shot general-purpose VLMs, specialized medical VLMs, and even larger-scale models.
-Despite its compact 2B parameter size, Med-R1 achieves competitive performance against 36×36$\times$36 × larger models, demonstrating that scalability does not have to come at the cost of efficiency.
-Its lightweight nature significantly lowers computational and deployment barriers, making it more practical for real-world clinical applications where resource constraints are critical.
-While VLM medical reasoning remains an emerging field with ongoing challenges, Med-R1 highlights the potential of reinforcement learning in developing safer, more transparent, and clinically reliable AI-driven healthcare solutions.
-Future research could further enhance Med-R1 by integrating additional multi-modal medical data, improving uncertainty quantification, and refining expert-aligned interpretability, ultimately advancing scalable and trustworthy medical reasoning systems.
-6 Acknowledgement
-This work is supported in part by the National Institutes of Health under award numbers R01CA272991, R01EB032680, R01DE033512, and U54CA274513.
+We present Med-R1, a RL-enhanced vision-language model that improves medical reasoning across diverse imaging modalities.
+By leveraging GRPO, Med-R1 surpasses the limitations of supervised fine-tuning, achieving strong cross-modality and cross-task generalization.
+Despite its compact 2B size, it matches or exceeds much larger models—including domain-specific and general-purpose VLMs—while remaining lightweight and deployable in resource-constrained clinical settings.
+Interestingly, we find that RL-based learning without explicit step-by-step rationales can outperform reasoning-heavy approaches, particularly under domain shift.
+This suggests that the quality, not quantity, of reasoning is key for robust medical adaptation.
+Med-R1 opens promising directions for developing scalable, reliable, and efficient AI systems in healthcare.

### v5

@@
 Abstract
-Vision-language models (VLMs) have achieved impressive progress in natural image reasoning, yet their potential in medical imaging remains underexplored.
-Medical vision-language tasks demand precise understanding and clinically coherent answers, which are difficult to achieve due to the complexity of medical data and the scarcity of high-quality expert annotations.
-These challenges limit the effectiveness of conventional supervised fine-tuning (SFT) and Chain-of-Thought (CoT) strategies that work well in general domains.
-To address these challenges, we propose Med-R1, a reinforcement learning (RL)-enhanced vision-language model designed to improve generalization and reliability in medical reasoning.
-Built on the DeepSeek strategy, Med-R1 adopts Group Relative Policy Optimization (GRPO) to encourage reward-guided learning beyond static annotations.
-We comprehensively evaluate Med-R1 across eight distinct medical imaging modalities.
-Med-R1 achieves a 29.94% improvement in average accuracy over its base model Qwen2-VL-2B, and even outperforms Qwen2-VL-72B—a model with 36$\times$ more parameters.
-To assess cross-task generalization, we further evaluate Med-R1 on five question types.
-Med-R1 outperforms Qwen2-VL-2B by 32.06% in question-type generalization, also surpassing Qwen2-VL-72B.
-We further explore the thinking process in Med-R1, a crucial component for the success of Deepseek-R1.
-Our results show that omitting intermediate rationales (No-Thinking-Med-R1) not only improves in-domain and cross-domain generalization with less training, but also challenges the assumption that more reasoning always helps.
-These findings suggest that in medical VQA, it is not reasoning itself, but its quality and domain alignment, that determine effectiveness.
-Together, these results highlight that RL improves medical reasoning and generalization, enabling efficient and reliable VLMs for real-world deployment.
+[…]
 {IEEEkeywords}
@@
 Unlike SFT, which aligns outputs to fixed supervision and often leads to shortcut learning, RL encourages exploration of diverse reasoning strategies through reward signals—without requiring explicit CoT annotations [11].
-We adopt Group Relative Policy Optimization (GRPO) [12], a lightweight alternative to PPO [13] that stabilizes training via rule-based rewards and group-relative comparisons.
+We adopt Group Relative Policy Optimization (GRPO) [12], a lightweight alternative to PPO (Proximal Policy Optimization) [13] that stabilizes training via rule-based rewards and group-relative comparisons.
 These mechanisms reduce computational overhead while encouraging clinically grounded reasoning, making GRPO well-suited for medical tasks where scalability and reliability are essential.
 Figure 1: Overview of the Evaluation Framework: Eight Medical Imaging Modalities and Five Medical Vision Question Answering Tasks.
-We evaluate Med-R1 across eight distinct medical imaging modalities—CT, MRI, Ultrasound, Dermoscopy, Fundus Photography, OCT, Microscopy Images, and X-ray Imaging—as well as five medical vision question answering tasks: anatomy identification, disease diagnosis, lesion grading, modality recognition, and biological attribute analysis.
+We evaluate Med-R1 across eight distinct medical imaging modalities—Computed Tomography (CT), Magnetic Resonance Imaging (MRI), Ultrasound, Dermoscopy, Fundus Photography, OCT (Optical coherence tomography ), Microscopy Images, and X-ray Imaging—as well as five medical vision question answering tasks: anatomy identification, disease diagnosis, lesion grading, modality recognition, and biological attribute analysis.
 Example images and corresponding clinical questions illustrate the diversity of medical data and the challenges in developing generalizable vision-language models for automated medical reasoning.
@@
 Med-R1 improves question-type generalization accuracy by 32.06% over Qwen2-VL-2B and surpasses Qwen2-VL-72B in this setting as well (Table 8).
-These demonstrate that RL enhances both parameter efficiency and task generalizability in medical VLMs.
-Finally, inspired by recent findings on reasoning behavior in rule-based RL for classification tasks [14], we further investigate the role of intermediate rationales in Med-R1.
-A widely held assumption in VLM development is that generating step-by-step reasoning (“Think”) enhances generalization and performance.
-However, our empirical analysis reveals that this may not hold in domain-specific settings like medical VQA.
-We compare two RL post-training strategies: one that guides the model to produce explicit rationales before answering, and another that directly outputs the final answer without any intermediate reasoning steps.
-Surprisingly, we observe that the latter (No-Thinking) strategy often yields better generalization across both imaging modalities and question types.
-We hypothesize that in the absence of high-quality CoT supervision, RL-trained models fall back on their pre-trained reasoning patterns, which are typically tuned for general-domain tasks.
-Under large domain shifts—such as differences in image characteristics, task goals, and clinical language—this can result in hallucinated or misaligned rationales that impair performance.
-These findings suggest that it is not the presence of reasoning per se, but rather its quality and domain alignment, that determine its utility in medical vision-language reasoning.
-Med-R1 is a generalizable medical VLM that generates interpretable reasoning explanations for medical VQA across eight imaging modalities (Figure 2). We summarized our contribution as follows:
+These demonstrate that RL enhances both parameter efficiency and generalizability in medical VLMs.
+Finally, we investigate the impact of intermediate rationales in Med-R1.
+Conventional wisdom suggests that explicit step-by-step reasoning (“Think”) enhances generalization, yet our results show this is not always true in medical VQA.
+We compare three RL post-training strategies: (1) Think, which generates rationales before answering; (2) No-Think, which directly predicts the answer; and (3) our proposed Think After, which performs concise reasoning after prediction.
+We find that No-Think improves generalization across modalities, while Think often causes hallucinated rationales due to domain shift.
+In contrast, Think After preserves interpretability without sacrificing accuracy, striking the best balance between reliability and explainability.
+These findings reveal that in specialized domains, the effectiveness of reasoning depends more on its quality, timing, and domain alignment than its length, challenging the notion that “more thinking is better.”
+Med-R1 a comprehensive and systematic study of rule-based RL for medical reasoning across eight imaging modalities (Figure 2). We summarized our contribution as follows:
   1. 1.
-    Multi-modality multi-task reasoning medical VLM.
-We propose Med-R1, the first reasoning medical VLM capable of supporting eight medical imaging modalities (CT, MRI, Ultrasound, etc.) across five unique tasks.
+    Multi-modality and multi-task medical reasoning VLM.
+We propose Med-R1, a comprehensive and systematic study of rule-based reinforcement learning for medical reasoning, supporting eight imaging modalities (CT, MRI, Ultrasound, etc.) across five distinct clinical tasks.
 We demonstrate that RL-based fine-tuning effectively promotes modality-specific as well as cross-modality reasoning in the medical domain without the need for token-level supervision as in SFT.
-The proposed Med-R1 is capable of generating step-by-step, accurate, and clinically plausible explanations.
+The proposed Med-R1 is capable of generating step-by-step, accurate, and plausible explanations.
   2. 2.
@@
   3. 3.
-    Rethinking the “More Thinking is Better” Assumption: Our results challenge the common belief that generating explicit thinking processes improves generalization.
-In medical VQA, we find that RL without reasoning steps yields more accurate predictions, while free-form reasoning—learned from general-domain data—can lead to hallucinations under domain shift.
-This highlights that in specialized domains—where image types, task goals, and language style differ significantly from general VLM training—the quality of reasoning, rather than its quantity, is critical for generalization.
+    Rethinking the “More Thinking is Better” Assumption: Our results challenge the common belief that generating longer or more explicit reasoning chains necessarily improves generalization.
+We find that reinforcement learning without explicit reasoning often yields higher accuracy, as free-form reasoning learned from general-domain data can induce hallucinations under domain shift.
+However, this may reduce reliability and interpretability in medical applications.
+To address this, we introduce Think-After—a reasoning scheme where the model provides rationalization for its chosen answer after the prediction.
+This design preserves interpretability while mitigating the instability introduced by lengthy reasoning chains, offering a balance between accuracy and explainability.
+Our findings suggest that the quality and timing of reasoning are critical for robust generalization.
 2 Related Works
@@
 Reinforcement learning offers an alternative by enabling emergent reasoning without explicit supervision.
-Concurrently, MedVLM-R1 [21] also applies GRPO-based RL to radiology VQA, which is conducted on a relatively small dataset without generalization analysis, making its reasoning robustness challenging to assess.
+Concurrently, MedVLM-R1 [20] also applies GRPO-based RL to radiology VQA, which is focuses on a single radiology-specific setting (training on MRI and testing on CT and X-ray) and reports generalization within radiology, using roughly 600 training samples.
 Nonetheless, it highlights the growing interest in RL for medical VLMs.
@@
 Specifically, we leverage the popular RL-based post-training method GRPO [22] to train a large base MLLM [23] across 8 different modalities for medical reasoning and compare it with zero-shot and SFT performance of popular existing VLMs.
-We introduce the details of the algorithm, our reward design, and data structure below.
+We introduce the details of the SFT and RL algorithm, our reward design, and data structure below.
+Supervised Fine-tuning: We performed supervised fine-tuning (SFT) on the Qwen2-VL-2B and Qwen2.5-VL-3 models.
+Specifically, we fine-tuned the model using mixed-precision training (bfloat16) with gradient accumulation and checkpointing enabled for memory efficiency.
+The maximum sequence length was set to 4096 tokens, allowing the model to process full visual-text reasoning sequences.
+Training was conducted for one epoch with a learning rate of $2\times 10^{-5}$, cosine learning-rate scheduling, and a warm-up ratio of 0.1.
+The per-device batch size was 4, and the gradient accumulation step was 16, resulting in an effective batch size of 64.
 3.1 Group Relative Policy Optimization (GRPO)
@@
 Darker cell shades indicate higher accuracy for corresponding training-test pairs in each column.
-The base model is Qwen2-VL-2B.
-RL fine-tuned VLM
-Train Test CT MRI X-Ray US Der FP OCT Micro Overall
-CT 98.30 74.43 80.61 51.93 64.17 65.39 70.52 66.13 71.44
-MRI 62.33 99.29 73.75 53.18 75.04 64.57 75.59 66.31 71.26
-X-Ray 79.67 68.29 95.85 49.32 70.75 76.87 70.28 67.75 72.35
-US 55.88 63.96 70.53 99.86 64.78 63.02 63.92 71.53 69.19
-Der 53.81 67.36 73.31 49.80 95.10 70.22 64.03 67.93 67.71
-FP 54.98 65.95 75.79 49.08 71.36 92.44 65.33 66.40 67.67
-OCT 64.48 70.80 77.09 52.65 72.89 73.41 99.17 66.85 72.17
-Micro 60.97 63.25 74.92 52.56 64.55 63.66 66.86 93.51 67.54
-Overall 66.30 71.67 77.73 57.31 72.33 71.20 71.96 70.80 69.91
+The base model is Qwen2.5-VL-3B.
+RL fine-tuned Qwen2.5-VL-3B
+Train Test CT MRI X-Ray Ultrasound Dermoscopy Fundus OCT Microscopy Overall
+CT $94.35\pm 0.79$ $70.85\pm 1.10$ $86.50\pm 1.70$ $35.44\pm 2.05$ $66.77\pm 2.53$ $76.41\pm 2.51$ $84.08\pm 2.48$ $66.58\pm 2.79$ $72.62\pm 0.65$
+MRI $72.79\pm 1.53$ $98.57\pm 0.29$ $72\pm 1.80 38.38$$38\pm 2.07$ $68.30\pm 2.53$ $79.78\pm 2.32$ $85.97\pm 2.36$ $67.12\pm 2.79$ $74.33\pm 0.64$
+X-Ray $78.90\pm 1.39$ $59.73\pm 1.20$ $93.50\pm 1.21$ $38.28\pm 2.10$ $65.85\pm 2.57$ $76.05\pm 2.55$ $78.89\pm 2.77$ $65.14\pm 2.79$ $69.54\pm 0.68$
+Ultrasound $62.20\pm 1.67$ $63.00\pm 1.19$ $79.57\pm 1.98$ $98.84\pm 0.46$ $64.09\pm 2.57$ $72.22\pm 2.60$ $78.54\pm 2.77$ $69.82\pm 2.75$ $73.53\pm 0.65$
+Dermoscopy $64.30\pm 1.65$ $60.39\pm 1.18$ $80.56\pm 1.92$ $40.21\pm 2.12$ $84.99\pm 1.91$ $71.86\pm 2.69$ $74.76\pm 2.89$ $65.23\pm 2.75$ $67.79\pm 0.70$
+Fundus $68.59\pm 1.59$ $61.32\pm 1.19$ $81.98\pm 1.86$ $37.42\pm 2.10$ $67.84\pm 2.49$ $90.62\pm 1.68$ $79.13\pm 2.71$ $66.76\pm 2.79$ $69.21\pm 0.67$
+OCT $81.06\pm 1.34$ $73.92\pm 1.06$ $83.65\pm 1.80$ $35.78\pm 2.03$ $68.99\pm 2.53$ $80.87\pm 2.32$ $98.70\pm 0.77$ $67.12\pm 2.79$ $73.76\pm 0.65$
+Microscopy $65.50\pm 1.62$ $62.34\pm 1.21$ $80.31\pm 1.95$ $37.37\pm 2.10$ $64.24\pm 2.60$ $1.40\pm 2.64 77$$7.36\pm 2.83$ $88.02\pm 1.94$ $68.32\pm 0.69$
+Overall $73.46\pm 1.51$ $68.77\pm 1.12$ $83.72\pm 1.79$ $45.21\pm 2.14$ $68.88\pm 2.45$ $77.40\pm 2.50$ $82.18\pm 2.53$ $69.47\pm 2.74$ $71.14\pm 0.67$
   • CT - Computed Tomography; MRI - Magnetic Resonance Imaging; US - Ultrasound; Der - Dermoscopy; FP - Fundus Photography.
   • OCT - Optical Coherence Tomography; Micro - Microscopy Images; X-Ray - X-Ray Imaging
-Table 2: Cross-Modality Generalization Performance of Med-R1 with RL Post-Training.
+Table 2: Cross-Modality Generalization Performance of Med-R1 with No-think RL Post-Training.
 Accuracy (%) across eight medical imaging modalities, where rows indicate training modalities and columns test modalities.
 Darker cell shades indicate higher accuracy for corresponding training-test pairs in each column.
-The base model is Qwen2-VL-2B.
-No-Think RL fine-tuned VLM
-Train Test CT MRI X-Ray US Der FP OCT Micro Overall
-CT 99.54 82.18 80.00 46.24 72.13 72.50 76.06 70.36 74.88
-MRI 79.27 99.72 74.74 42.86 79.94 73.50 84.67 69.10 75.48
-X-Ray 78.37 69.94 97.09 46.43 70.75 80.97 69.81 70.63 73.00
-US 59.82 66.53 73.37 100.00 70.60 68.03 68.63 79.28 73.28
-Der 56.34 67.94 73.87 46.14 98.00 69.67 67.45 67.57 68.37
-FP 59.55 68.62 73.93 51.59 72.43 93.99 72.64 70.36 70.39
-OCT 80.47 73.50 77.15 49.08 65.08 70.95 100.00 67.21 72.93
-Micro 64.15 66.56 76.84 53.13 68.91 70.95 73.47 98.56 71.50
-Overall 72.19 74.37 78.37 54.43 74.73 75.07 76.59 74.13 72.48
+The base model is Qwen2.5-VL-3B
+No-Think RL fine-tuned Qwen2.5-VL-3B
+Train Test CT MRI X-Ray Ultrasound Dermoscopy Fundus OCT Microscopy Overall
+CT $98.27\pm 0.45$ $74.77\pm 1.04$ $85.33\pm 1.70$ $37.61\pm 2.05$ $69.07\pm 2.53$ $78.87\pm 2.46$ $87.26\pm 2.30$ $67.75\pm 2.75$ $74.87\pm 0.64$
+MRI $69.70\pm 1.56$ $99.61\pm 0.15$ $81.73\pm 1.89$ $34.91\pm 2.03$ $75.19\pm 2.34$ $80.51\pm 2.32$ $89.98\pm 2.00$ $67.48\pm 2.75$ $74.89\pm 0.65$
+X-Ray $80.10\pm 1.36$ $69.78\pm 1.14$ $95.79\pm 0.99$ $34.33\pm 2.05$ $71.29\pm 2.41$ $81.79\pm 2.28$ $83.14\pm 2.54$ $68.65\pm 2.79$ $73.11\pm 0.63$
+Ultrasound $56.40\pm 1.70$ $64.93\pm 1.18$ $80.25\pm 1.92$ $100.0\pm 0.00$ $68.38\pm 2.53$ $75.32\pm 2.60$ $76.30\pm 2.89$ $71.71\pm 2.66$ $74.16\pm 0.66$
+Dermoscopy $55.45\pm 1.70$ $66.48\pm 1.15$ $80.37\pm 1.95$ $37.42\pm 2.12$ $92.88\pm 1.38$ $71.86\pm 2.69$ $76.53\pm 2.83$ $67.12\pm 2.70$ $68.51\pm 0.69$
+Fundus $58.65\pm 1.68$ $66.08\pm 1.16$ $82.79\pm 1.83$ $34.33\pm 2.10$ $68.68\pm 2.49$ $93.81\pm 1.41$ $79.72\pm 2.71$ $65.86\pm 2.79$ $68.74\pm 0.69$
+OCT $80.38\pm 1.36$ $75.02\pm 1.05$ $86.81\pm 1.67$ $36.11\pm 2.05$ $72.21\pm 2.41$ $81.60\pm 2.28$ $99.88\pm 0.18$ $65.32\pm 2.79$ $74.67\pm 0.64$
+Microscopy $0.17\pm 1.73 66$$6.59\pm 1.16$ $83.10\pm 1.86$ $36.60\pm 2.10$ $1.29\pm 2.45 77$$7.05\pm 2.46$ $80.66\pm 2.71$ $97.66\pm 0.86$ $71.64\pm 0.67$
+Overall $69.89\pm 1.59$ $72.91\pm 1.08$ $84.52\pm 1.73$ $43.91\pm 2.14$ $73.62\pm 2.37$ $80.10\pm 2.41$ $84.18\pm 2.47$ $71.44\pm 2.65$ $72.57\pm 0.66$
+Table 3: Cross-Modality Generalization Performance of Med-R1 with Think-after RL Post-Training.
+Accuracy (%) across eight medical imaging modalities, where rows indicate training modalities and columns test modalities.
+Darker cell shades indicate higher accuracy for corresponding training-test pairs in each column.
+The base model is Qwen2.5-VL-3B.
+Think-after RL fine-tuned Qwen2.5-VL-3B
+Train Test CT MRI X-Ray Ultrasound Dermoscopy Fundus OCT Microscopy Overall
+CT $96.64\pm 0.62$ $72.53\pm 1.08$ $85.45\pm 1.67$ $36.31\pm 2.03$ $66.16\pm 2.53$ $79.78\pm 2.37$ $86.20\pm 2.30$ $68.20\pm 2.75$ $73.91\pm 0.65$
+MRI $63.65\pm 1.67$ $98.87\pm 0.26$ $81.11\pm 1.92$ $33.70\pm 2.03$ $71.82\pm 2.45$ $81.51\pm 2.32$ $86.91\pm 2.30$ $67.57\pm 2.70$ $73.14\pm 0.67$
+X-Ray $79.20\pm 1.42$ $72.70\pm 1.10$ $94.61\pm 1.08$ $31.58\pm 2.00$ $70.21\pm 2.45$ $80.60\pm 2.32$ $85.97\pm 2.30$ $68.65\pm 2.70$ $72.94\pm 0.66$
+Ultrasound $59.06\pm 1.70$ $66.23\pm 1.15$ $80.25\pm 1.95$ $99.37\pm 0.34$ $67.38\pm 2.53$ $74.95\pm 2.55$ $78.77\pm 2.77$ $67.93\pm 2.75$ $74.24\pm 0.64$
+Dermoscopy $56.77\pm 1.71$ $66.86\pm 1.15$ $80.37\pm 1.95$ $34.72\pm 2.07$ $89.51\pm 1.68$ $75.05\pm 2.55$ $78.66\pm 2.77$ $67.30\pm 2.75$ $68.65\pm 0.67$
+Fundus $58.01\pm 1.68$ $65.89\pm 1.16$ $81.30\pm 1.86$ $32.84\pm 2.03$ $68.68\pm 2.53$ $92.53\pm 1.55$ $79.36\pm 2.77$ $66.22\pm 2.79$ $68.10\pm 0.69$
+OCT $79.98\pm 1.37$ $75.45\pm 1.05$ $85.51\pm 1.73$ $34.14\pm 2.03$ $71.21\pm 2.45$ $81.69\pm 2.32$ $99.29\pm 0.53$ $66.94\pm 2.75$ $74.28\pm 0.65$
+Microscopy $59.27\pm 1.68$ $64.22\pm 1.19$ $81.11\pm 1.89$ $34.47\pm 2.05$ $63.86\pm 2.60$ $73.68\pm 2.64$ $78.42\pm 2.77$ $86.67\pm 2.03$ $67.71\pm 0.69$
+Overall $69.07\pm 1.62$ $72.84\pm 1.10$ $83.72\pm 1.76$ $42.14\pm 2.12$ $71.10\pm 2.53$ $79.97\pm 2.37$ $84.20\pm 2.47$ $69.93\pm 2.65$ $71.62\pm 0.66$
+  • CT - Computed Tomography; MRI - Magnetic Resonance Imaging; US - Ultrasound; Der - Dermoscopy; FP - Fundus Photography.
+  • OCT - Optical Coherence Tomography; Micro - Microscopy Images; X-Ray - X-Ray Imaging
 Reward design: We follow [12] and use two types of reward: format and accuracy.
@@
 In practice, the ground truths are letter options “A, B, C, D" for multiple choice questions, and we treat all responses with correct letter options as the leading word as correct (“A…").
-3.2 No-Thinking-Med-R1
+3.2 No-Thinking Med-R1
 Previous work [26] found that, in image classification tasks, removing both reasoning and format supervision during RL could sometimes improve performance.
@@
 Modality abbreviations: CT – Computed Tomography; MRI – Magnetic Resonance Imaging; US – Ultrasound; Der – Dermoscopy; FP – Fundus Photography; OCT – Optical Coherence Tomography; Micro – Microscopy; X-Ray – X-ray Imaging.
-Methods Modality CT MRI X-Ray US Der FP OCT Micro Overall
+Methods Modality CT MRI X-Ray Ultrasound Dermoscopy Fundus OCT Microscopy Overall
 Zero-shot VLMs
-[…]
+$\text{BLIP-2}^{\dagger}$ [15] 56.74 41.32 67.58 37.27 40.65 46.24 68.08 50.40 51.04
 $\text{InstructBLIP}^{\dagger}$ [27] 28.72 33.15 61.04 41.25 62.22 50.31 42.59 46.29 45.70
 $\text{LLaVA}^{\dagger}$ [28] 17.73 26.72 30.70 18.66 49.74 47.11 33.73 28.87 31.66
-[…]
-[…]
+$\text{LLaMA Adapter v2}^{\dagger}$ [29] 21.41 26.63 46.44 34.05 51.76 50.74 33.00 38.66 37.83
+$\text{MiniGPT-4}^{\dagger}$ [30] 22.81 27.48 38.30 25.50 40.25 38.33 31.40 36.23 32.54
+InternVL2 [31] 40.20 58.10 57.90 49.10 51.90 53.20 59.10 64.00 54.19
 Qwen2-VL-2B [3] 45.10 38.57 39.32 30.86 35.83 43.17 35.14 36.85 38.11
@@
 Zero-shot Medical VLMs
-[…]
+$\text{LLaVA-Med}^{\dagger}$ [8] 18.69 27.47 30.68 29.88 44.95 39.03 34.61 33.29 32.33
 $\text{RadFM}^{\dagger}$ [34] 27.56 24.06 30.95 16.57 39.21 36.89 32.80 27.97 29.50
-[…]
+$\text{Med-Flamingo}^{\dagger}$ [6] 38.47 40.56 30.34 24.64 32.43 30.12 26.51 19.93 30.38
 $\text{MedVInT}^{\dagger}$ [7] 40.74 43.10 55.10 41.26 29.11 31.84 23.26 32.00 37.05
+HuatuoGPT-Vision [9] 35.30 40.40 41.50 60.10 53.10 51.40 59.30 62.30 50.43
+HealthGPT [35] 35.50 78.50 81.90 51.40 64.90 54.60 89.30 88.20 68.04
 Fine-tuned VLMs
@@
 Qwen2.5-VL-3B (SFT) 56.06 60.81 69.23 41.77 60.11 69.19 63.95 65.66 60.85
-Qwen2-VL-2B (GRPO + Think) 66.30 71.67 77.73 57.31 72.33 71.20 71.96 70.80 69.91
-Qwen2-VL-2B (GRPO + Nothink) 72.19 74.37 78.37 54.43 74.73 75.07 76.59 74.13 72.49
-Qwen2.5-VL-3B (GRPO + Nothink) 69.89 72.91 84.52 43.91 73.62 80.10 84.18 71.44 72.57
+Qwen2-VL-2B (Think) 66.30 71.67 77.73 57.31 72.33 71.20 71.96 70.80 69.91
+Qwen2-VL-2B (Nothink) 72.19 74.37 78.37 54.43 74.73 75.07 76.59 74.13 72.49
+Qwen2.5-VL-3B (Think) 73.46 68.77 83.72 45.21 68.88 77.40 82.18 69.47 71.14
+Qwen2.5-VL-3B (Think-after) 69.07 72.84 83.72 42.14 71.10 79.97 84.20 69.93 71.62
+Qwen2.5-VL-3B (Nothink) 69.89 72.91 84.52 43.91 73.62 80.10 84.18 71.44 72.57
+3.3 Think-after Med-R1
+As said in the previous section, removing both reasoning and format supervision during RL could sometimes improve performance.
+However, as medical AI systems must not only achieve high accuracy but also provide reasoning that physicians can review and validate.
+To address this, we introduce a new reasoning protocol termed Think-After, in which the model first predicts the answer and then generates a post-hoc rationale explaining that decision.
+This design preserves interpretability while minimizing the instability introduced by lengthy reasoning chains.
+Concretely, the instruction prompt is revised as: {Question}.
+Output the single-letter choice (A, B, C, D,...) in <answer>...</answer> tags.
+Then provide the reasoning step by step to explain why you chose this answer.
+where {Question} is replaced by each specific VQA item.
+By separating answer prediction and rationale generation, we reduce interference between reasoning and decision processes, achieving balanced trade-off between accuracy and interpretability.
 4 Experiment & Results
@@
 We adopt the VQA data from the open-access part of the OmniMedVQA benchmark [36], which consists of a total of $82,059$ images and $88,996$ vision question answering pairs.
-OmniMedVQA includes VQA pairs from eight imaging modalities: CT ($15,808$), MRI ($31,877$), X-Ray ($7,916$), Ultrasound ($10,991$), Dermoscopy ($6,679$), Fundus ($5,398$), OCT ($,64$ 646), and Microscopy ($5,680$).
+OmniMedVQA includes VQA pairs from eight imaging modalities: CT ($15,808$), MRI ($31,877$), X-Ray ($7,916$), Ultrasound ($10,991$), Dermoscopy ($6,679$), Fundus ($5,398$), OCT ($4,646$), and Microscopy ($5,680$).
 It is also categorized into five VQA question types, including Anatomy Identification ($16,448$), Disease Diagnosis ($55,387$), Lesion Grading ($2,098$), Modality Recognition ($11,565$), and Other Biological Attributes ($3,498$).
@@
   • Cross-task generalization: We identify 5 distinct task types within the dataset and adopt the same train-test partitioning strategy as in the cross-modality setting, training on one task and evaluating on the other four.
+Table 5: Cross-Task Generalization of Med-R1: Performance is evaluated across five clinical reasoning task types (rows: training tasks, columns: test tasks), with darker cell shading indicating stronger generalization.
+Results demonstrate that domain-specific training (e.g., Disease Diagnosis) preserves in-task expertise while maintaining adaptability to unseen tasks, particularly for modality-agnostic skills like Modality Recognition.
+The base model is Qwen2.5-VL-3B.
+RL fine-tuned Qwen2.5-VL-3B
+Train Test Anatomy Disease Lesion Modality Other Overall
+Identification Diagnosis Grading Recognition Attributes
+Anatomy Identification $94.65\pm 0.75$ $61.87\pm 0.90$ $63.30\pm 4.59$ $96.66\pm 0.73$ $78.69\pm 3.05$ $79.04\pm 0.65$
+Disease Diagnosis $43.83\pm 1.67$ $97.74\pm 0.28$ $79.82\pm 3.67$ $96.66\pm 0.73$ $87.78\pm 2.41$ $81.16\pm 0.49$
+Lesion Grading $47.10\pm 1.71$ $59.33\pm 0.93$ $88.99\pm 2.98$ $96.03\pm 0.77$ $69.03\pm 3.41$ $72.10\pm 0.70$
+Modality Recognition $45.96\pm 1.67$ $58.63\pm 0.91$ $67.89\pm 4.36$ $99.08\pm 0.40$ $69.46\pm 3.41$ $68.20\pm 0.71$
+Other Attributes $44.04\pm 1.68$ $60.09\pm 0.92$ $60.09\pm 4.59$ $95.74\pm 0.82$ $94.74\pm 1.63$ $70.94\pm 0.70$
+Overall $55.12\pm 0.75$ $67.53\pm 0.39$ $72.02\pm 1.86$ $96.83\pm 0.31$ $79.94\pm 1.32$ $74.29\pm 0.65$
+Table 6: No-Think Cross-Task Generalization of Med-R1: Performance is evaluated across five clinical reasoning task types (rows: training tasks, columns: test tasks), with darker cell shading indicating stronger generalization.
+No-Think RL fine-tuned Qwen2.5-VL-3B
+Train Test Anatomy Disease Lesion Modality Other Overall
+Identification Diagnosis Grading Recognition Attributes
+Anatomy Identification $96.85\pm 0.57$ $63.86\pm 0.90$ $79.59\pm 3.78$ $96.82\pm 0.69$ $77.41\pm 3.12$ $82.91\pm 0.64$
+Disease Diagnosis $44.82\pm 1.73$ $98.78\pm 0.21$ $85.78\pm 3.33$ $95.82\pm 0.82$ $90.91\pm 2.13$ $83.22\pm 0.48$
+Lesion Grading $44.61\pm 1.70$ $60.27\pm 0.91$ $97.02\pm 1.61$ $95.90\pm 0.79$ $82.10\pm 2.84$ $75.98\pm 0.70$
+Modality Recognition $47.07\pm 1.71$ $59.58\pm 0.91$ $75.92\pm 3.90$ $99.75\pm 0.19$ $81.53\pm 2.84$ $72.77\pm 0.70$
+Other Attributes $43.98\pm 1.71$ $61.11\pm 0.90$ $61.70\pm 4.59$ $96.07\pm 0.77$ $95.60\pm 1.49$ $71.69\pm 0.71$
+Overall $55.46\pm 0.77$ $68.72\pm 0.39$ $80.00\pm 1.67$ $96.87\pm 0.31$ $85.51\pm 1.16$ $77.31\pm 0.62$
+We focus on VQA, which integrates core vision–language abilities such as classification, grounding, and reasoning into a unified framework, making it a representative precursor to detection or captioning tasks.
 Baseline Methods & Evaluation Metric.
-Our evaluation framework integrates three classes of baselines: general VLMs including BLIP-2 [16], InstructBLIP [27], and the Qwen2-VL series (2B/7B/72B) [3, 30]; medical VLMs pretrained on domain-specific corpora (Med-Flamingo [6], MedVInT [7], LLaVA-Med [8]); and SFT variants of Qwen2-VL-2B.
-We evaluate model performance using VQA choice accuracy, the standard metric for medical VQA, where models select answers from clinically validated options.
-Given a medical image $I$, a question $Q$, and $K$ candidate answers $\{A_{k}\}_{k=1}^{K}$ (one correct and $K-1$ distractors), accuracy is computed as:
-$\text{Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left(\hat{y}_{i}=y_{i}\right)$ (2)
-where $N$ is the total number of test cases, $\hat{y}_{i}$ is the model’s predicted answer index, $y_{i}$ is the ground truth index, and $\mathbb{I}$ is the indicator function, which returns 1 if the predicted answer matches the ground truth and 0 otherwise.
+We report our results by separating the baselines into three groups.
+Zero-shot VLMs are models pre-trained for general-purpose VQA without medical adaptation.
+Medical VLMs are models pre-trained on medical data specifically for medical VQA.
+Fine-tuned VLMs are models trained on OmniMedVQA using supervised fine-tuning and RL fine-tuning.
+We evaluate model performance using VQA choice accuracy, the standard metric for medical VQA, where the model selects the correct answer from $K$ clinically validated options.
+Given an image $I$, a question $Q$, and candidate answers $\{A_{k}\}_{k=1}^{K}$, accuracy is defined as:
+$\text{Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}(\hat{y}_{i}=y_{i}),$ (2)
+where $N$ is the total number of test cases, $\hat{y}_{i}$ is the predicted answer index, $y_{i}$ the ground truth, and $\mathbb{I}$ the indicator function (1 if correct, 0 otherwise).
+To assess result reliability, we computed 95% bootstrap confidence intervals (CIs) for all model variants across modalities and tasks.
+Each experiment performed 10,000 bootstrap resamplings of binary accuracy labels to estimate the mean and confidence bounds, reporting $mean\pm half\text{-}width$ for each model.
+All fine-tuned VLM results (SFT, GRPO+Think, GRPO+Think-After, and GRPO+No-Think) are computed under the identical averaging protocol—macro-averaged accuracy across all cross-modality training–testing pairs—to ensure fair and consistent comparison.
 4.2 Cross-Modality Generalization
@@
 Our experiments focus on two key aspects: (1) cross-modal generalization, where the model is trained on one modality and tested on another, and (2) comparative performance against other popular VLMs and medical-specific VLMs evaluated using zero-shot and SFT.
-Table 4: Cross-Task Generalization of Med-R1: Performance is evaluated across five clinical reasoning task types (rows: training tasks, columns: test tasks), with darker cell shading indicating stronger generalization.
-Results demonstrate that domain-specific training (e.g., Disease Diagnosis) preserves in-task expertise while maintaining adaptability to unseen tasks, particularly for modality-agnostic skills like Modality Recognition.
-The base model is Qwen2-VL-2B.
+Table 7: Think-After Cross-Task Generalization of Med-R1: Performance is evaluated across five clinical reasoning task types (rows: training tasks, columns: test tasks), with darker cell shading indicating stronger generalization.
+Think-After fine-tuned Qwen2.5-VL-3B
 Train Test Anatomy Disease Lesion Modality Other Overall
 Identification Diagnosis Grading Recognition Attributes
-Anatomy Identification 96.06 59.07 55.51 98.62 74.86 76.83
-Disease Diagnosis 54.16 98.25 73.85 97.87 84.09 81.64
-Lesion Grading 52.30 57.72 86.24 97.32 74.29 73.57
-Modality Recognition 55.30 55.43 56.88 99.46 70.88 67.59
-Other Attributes 56.60 59.93 56.88 97.91 96.59 73.58
-Overall 62.88 66.08 65.87 98.24 80.14 74.64
-Table 5: No-Think Cross-Task Generalization of Med-R1: Performance is evaluated across five clinical reasoning task types (rows: training tasks, columns: test tasks), with darker cell shading indicating stronger generalization.
-The base model is Qwen2-VL-2B.
-Train Test Anatomy Disease Lesion Modality Other Overall
-Identification Diagnosis Grading Recognition Attributes
-Anatomy Identification 96.33 59.29 64.22 99.04 76.56 79.09
-Disease Diagnosis 55.00 98.30 63.76 98.08 85.23 80.07
-Lesion Grading 54.22 58.04 96.33 97.37 72.30 75.65
-Modality Recognition 56.84 57.95 55.05 99.79 73.15 68.56
-Other Attributes 56.29 58.00 52.29 97.90 99.29 72.75
-Overall 63.74 66.32 66.33 98.44 81.31 75.22
+Anatomy Identification $95.22\pm 0.74$ $60.62\pm 0.90$ $65.83\pm 4.36$ $96.57\pm 0.73$ $81.53\pm 2.91$ $79.95\pm 0.65$
+Disease Diagnosis $43.11\pm 1.68$ $97.86\pm 0.27$ $81.65\pm 3.67$ $95.99\pm 0.77$ $92.33\pm 1.99$ $82.19\pm 0.49$
+Lesion Grading $43.56\pm 1.70$ $59.78\pm 0.92$ $85.55\pm 3.21$ $96.20\pm 0.77$ $78.55\pm 3.05$ $72.73\pm 0.71$
+Modality Recognition $44.73\pm 1.71$ $58.82\pm 0.93$ $68.12\pm 4.47$ $98.87\pm 0.44$ $78.12\pm 3.05$ $69.73\pm 0.71$
+Other Attributes $45.21\pm 1.67$ $59.65\pm 0.90$ $67.43\pm 4.36$ $95.74\pm 0.82$ $95.74\pm 1.49$ $72.75\pm 0.71$
+Overall $54.36\pm 0.75$ $67.34\pm 0.40$ $73.72\pm 1.86$ $96.67\pm 0.32$ $85.26\pm 1.18$ $75.47\pm 0.63$
 Results on generalization.
@@
 Zero-shot VLMs
-[…]
+$\text{BLIP-2}^{\dagger}$ [15] 44.39 44.51 29.03 68.19 67.95 48.12
 $\text{InstructBLIP}^{\dagger}$ [27] 44.35 32.29 59.25 75.27 23.72 40.40
 $\text{LLaVA}^{\dagger}$ [28] 25.86 29.10 43.95 21.36 31.90 27.96
-[…]
-[…]
+$\text{LLaMA Adapter v2}^{\dagger}$ [29] 33.72 31.19 41.99 37.29 34.22 32.82
+$\text{MiniGPT-4}^{\dagger}$ [30] 28.88 30.47 34.56 26.43 30.36 29.74
 Qwen2-VL-2B [3] 30.70 36.53 43.58 59.90 42.19 42.58
@@
 Zero-shot Medical VLMs
-[…]
+$\text{LLaVA-Med}^{\dagger}$ [8] 29.53 29.22 34.18 26.93 33.08 29.25
 $\text{RadFM}^{\dagger}$ [34] 13.31 21.69 30.35 26.64 43.85 26.99
-[…]
+$\text{Med-Flamingo}^{\dagger}$ [6] 24.93 38.90 30.74 30.19 14.18 34.03
 $\text{MedVInT}^{\dagger}$ [7] 40.26 35.78 12.77 68.10 30.30 40.04
@@
 Qwen2.5-VL-3B (SFT) 54.91 57.75 64.04 84.95 71.56 66.64
-Qwen2-VL-2B (GRPO + Think) 62.88 66.08 65.87 98.24 80.14 74.64
-Qwen2-VL-2B (GRPO + Nothink) 63.74 66.32 66.33 98.44 81.31 75.22
-Qwen2.5-VL-3B (GRPO + Nothink) 55.47 68.72 80.00 96.87 85.51 77.31
+Qwen2-VL-2B (Think) 62.88 66.08 65.87 98.24 80.14 74.64
+Qwen2-VL-2B (Nothink) 63.74 66.32 66.33 98.44 81.31 75.22
+Qwen2.5-VL-3B (Think) 55.12 67.53 72.02 96.83 79.94 74.29
+Qwen2.5-VL-3B (Think-after) 54.36 67.34 73.72 96.67 85.26 75.47
+Qwen2.5-VL-3B (Nothink) 55.47 68.72 80.00 96.87 85.51 77.31
 Comparison to zero-shot and SFT evaluations with other VLMs.
@@
 For zero-shot results, each cell denotes the zero-shot evaluation accuracy of the model on the particular modality.
-For the fine-tuned VLM results (last two rows), each cell represents the overall accuracy, reflecting the average generalization performance of the given modality when evaluated using models that were separately fine-tuned on each of the eight training modalities.
-Against general-purpose VLMs, our 2B-parameter model achieves 69.91% overall accuracy, surpassing the 72B-parameter Qwen2-VL by 1.86%—a notable result given the 36×36$\times$36 × parameter disparity.
+For all the fine-tuned VLM results, each cell represents the overall accuracy, reflecting the average generalization performance of the given modality when evaluated using models that were separately fine-tuned on each of the eight training modalities.
+Against general-purpose VLMs, our 2B-parameter model achieves 69.91% overall accuracy, surpassing the 72B-parameter Qwen2-VL by 1.86%—a notable result given the $36\times$ parameter disparity.
 This advantage amplifies in critical diagnostic tasks: Med-R1 attains 71.67% accuracy in MRI compared to Qwen2-VL-72B’s 69.39%, and achieves 72.33% versus 65.31% in dermoscopy, challenging the prevailing scale-equals-performance paradigm.
@@
 53.99%).
+Figure 3: Training accuracy–reward curves of the Think, No-Think, and Think-After models on CT and MRI datasets.
 4.3 Cross-Task Generalization
@@
 Each cell shows the performance gap (%) when trained on one modality (y-axis) and evaluated on others (x-axis).
-Red indicates improvement; blue indicates degradation.
-Figure 4: Cross-task accuracy difference between No-Thinking-Med-R1 and Med-R1.
-Each cell shows the performance gap (%) when trained on one task (y-axis) and tested on another (x-axis).
 Red indicates improvement; blue indicates degradation.
@@
 These results highlight the trade-off between specialization and adaptability, emphasizing the importance of task selection when designing models for broad medical applications.
+Figure 5: Cross-task accuracy difference between No-Thinking-Med-R1 and Med-R1.
+Each cell shows the performance gap (%) when trained on one task (y-axis) and tested on another (x-axis).
+Red indicates improvement; blue indicates degradation.
 Comparison to zero-shot and SFT evaluations with other VLMs.
@@
 In contrast, the average generalization with the identical base model trained with SFT is merely 63.39%, 11.25% below Med-R1, further demonstrating the strong generalization capability of Med-R1.
-Figure 5: Rewards & Length.
-The top row shows the change in RL rewards during Med-R1 training across eight imaging modalities and five task types.
-We observe that the model quickly converges to stable reward values within 100–200 steps, indicating the computational efficiency of our RL post-training process.
-The bottom row shows the output completion length over training steps.
-Interestingly, the output length gradually decreases as training progresses, suggesting that the reward function discourages overly verbose or unfaithful reasoning.
-This observation challenges the conventional assumption that longer reasoning necessarily benefits vision-language performance in VQA.
-4.4 Analysis of Thinking Process
+4.4 Analysis of No-Think
 We present the results of No-Thinking-Med-R1 in this section.
@@
 Med-R1 demonstrates that even without reasoning supervision, RL can offer a robust and efficient path for adapting VLMs to medicine.
-4.5 Analysis of Rewards and Completion Length.
-The top row of Figure 5 shows the reward dynamics during GRPO-based reinforcement learning.
-Across all modalities and tasks, Med-R1 converges within 100–200 steps, as evidenced by the log-scale reward curves.
-This rapid and consistent convergence highlights the efficiency of our training pipeline and the model’s ability to quickly adapt to the reward structure.
-Moreover, the smooth reward trajectories suggest that GRPO remains stable even with limited RL updates, making it well-suited for low-resource scenarios.
-Unlike in scientific domains such as mathematics [12], where model responses become increasingly detailed as thinking deepens, we observe a different trend in Med-RL training: the average response length decreases rapidly and then fluctuates (Figure 5).
-Rather than generating elaborate rationales, the model learns to produce shorter, more direct outputs—potentially favoring conciseness over verbosity.
-This aligns with our findings in subsection 4.4, suggesting that extended reasoning may not always benefit cross-domain VQA, especially under large domain shifts.
-We hypothesize that RL without CoT supervision encourages the model to rely on its pre-trained knowledge when generating rationales.
-While this may be effective in general domains, it can lead to hallucinated or misaligned reasoning in specialized domains like medical images.
-These results challenge the common assumption that more thinking always improves performance, and highlight that the “quality”—rather than the “presence”—of thinking is a crucial factor affecting the outcome of RL post-training.
+4.5 Analysis of Think-after
+We introduce the Think-After strategy primarily to address the practical need for achieving both high accuracy and interpretable reasoning in medical VLMs (Table 7, Table 3).
+Rather than serving solely as a control experiment, Think-After is designed to meet the dual objective of preserving predictive performance while providing reasoning traces that clinicians can review and validate.
+Beyond satisfying the need for both accuracy and interpretability, Think-After also provides insight into why the No-Thinking strategy sometimes outperforms the conventional Thinking approach.
+As shown in Figure 3, Think-After achieves faster convergence and higher accuracy rewards than the Thinking variant, suggesting that generating reasoning tokens before the answer may disrupt the autoregressive generation process and hinder optimization.
+However, Think-After still performs slightly below No-Thinking, indicating that additional mechanisms—such as residual contextual coupling or reasoning-token noise—may also contribute to the performance gap.
+While Think-After does not completely resolve the question of why No-Thinking surpasses both pre- and post-answer reasoning, it fulfills the essential requirement of combining strong accuracy with interpretable outputs, and simultaneously offers valuable clues toward understanding the dynamics between reasoning generation and performance in medical VLMs.
+Table 9: Reader Study Results. Three readers independently evaluated 100 correctly answered VQA samples for factual correctness (Quality, 1–5) and reasoning–answer consistency (Consistency, %).
+Reader Quality Consistency (%)
+A 4.23 93.0
+B 4.19 92.0
+C 4.34 95.0
+Mean 4.25 93.3
+  • Quality reflects factual and clinical correctness of model answers; Consistency measures logical alignment between reasoning and answers.
+4.6 Reader Study
+To further assess interpretability and clinical relevance, we conducted a reader study with three researchers experienced in medical imaging and vision–language models.
+Each independently evaluated 100 representative VQA samples generated by different model variants, covering diverse imaging modalities and reasoning types (Table 9).
+Evaluations focused on (1) the factual and clinical correctness of model answers and (2) the logical consistency between the reasoning (“Think”) and the predicted answer (“Answer”).
+Across readers, the Think-After model achieved the highest agreement and was consistently judged to produce reasoning that was both coherent and clinically sound, demonstrating improved interpretability and reasoning reliability for medical decision-support applications.
 4.7 Limitations and Future Work.
-This work marks an initial step in applying RL to medical vision-language models.
+This work marks an initial step in applying RL to medical vision–language models.
 We adopt a frame-level VQA setting for consistent evaluation across modalities, but this simplifies the real-world complexity of medical imaging.
 In practice, CT and MRI are volumetric, and ultrasound is dynamic, requiring reasoning across slices and time.
-Future directions include extending Med-R1 to support multiple inputs.
-We also plan to incorporate patient context and explore more advanced reasoning frameworks for clinical deployment.
+Another limitation lies in the reasoning supervision and data scale.
+Our findings suggest that the “No-Thinking” model occasionally outperforms reasoning-enabled variants, which may stem from the limited availability of high-quality medical reasoning data.
+In domains where clinically faithful chain-of-thought (CoT) annotations are scarce, the model may not learn to benefit from explicit reasoning, leading the reasoning process to introduce noise rather than insight.
+We believe this reflects a broader limitation of current medical datasets rather than of reasoning itself.
+Future work should explore scaling up medically grounded CoT data and aligning RL rewards with clinically validated reasoning quality to fully realize the benefits of explicit reasoning.
+Future directions also include extending Med-R1 to support multi-frame or volumetric inputs, incorporating patient context, and investigating more advanced reasoning frameworks for clinical deployment.
 5 Conclusion
-We present Med-R1, a RL-enhanced vision-language model that improves medical reasoning across diverse imaging modalities.
-By leveraging GRPO, Med-R1 surpasses the limitations of supervised fine-tuning, achieving strong cross-modality and cross-task generalization.
-Despite its compact 2B size, it matches or exceeds much larger models—including domain-specific and general-purpose VLMs—while remaining lightweight and deployable in resource-constrained clinical settings.
-Interestingly, we find that RL-based learning without explicit step-by-step rationales can outperform reasoning-heavy approaches, particularly under domain shift.
-This suggests that the quality, not quantity, of reasoning is key for robust medical adaptation.
-Med-R1 opens promising directions for developing scalable, reliable, and efficient AI systems in healthcare.
+We present Med-R1, a reinforcement learning–enhanced vision–language model for improving medical reasoning across diverse imaging modalities and clinical tasks.
+Leveraging GRPO-based post-training, Med-R1 achieves strong cross-modality and cross-task generalization, surpassing the limits of supervised fine-tuning.
+Despite its compact 3B scale, Med-R1 performs competitively with or better than larger medical and general VLMs, while remaining efficient for deployment.
+Through systematic analysis, we find that removing explicit reasoning (No-Think) improves convergence and generalization, suggesting that reasoning quality, not quantity, drives medical performance.
+To balance interpretability and accuracy, we introduce the Think-After strategy, which decouples reasoning from answer generation and enhances clinical transparency without compromising accuracy.
+Overall, Med-R1 establishes a scalable framework for reinforcement learning in medical VLMs and offers new insights into how reasoning dynamics interact with domain generalization, paving the way toward reliable and interpretable medical AI systems.